### v1

# General limit theorems for mixtures of free, monotone, and boolean independence

Abstract.
We study mixtures of free, monotone, and boolean independence described by directed graphs (digraphs). For a sequence of digraphs $G_{n}=(V_{n},E_{n})$, we give sufficient conditions for the limit $\widehat{\mu}=\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})$ to exist whenever the boolean convolution powers $\mu_{n}^{\uplus|V_{n}|}$ converge to some $\mu$. This in particular includes central limit and Poisson limit theorems, as well as limit theorems for each classical domain of attraction. The hypothesis on the sequence of $G_{n}$ is that the normalized counts of digraph homomorphisms from rooted trees into $G_{n}$ converge as $n\to\infty$, and we verify this for several families of examples where the $G_{n}$’s converge in some sense to a continuum limit. In particular, we obtain a new limit theorems for multiregular digraphs, as well as recovering several limit theorems in prior work.
Key words and phrases:
Noncommutative probability, limit theorem, digraph, free independence, monotone independence, boolean independence, bm-independence, tree, non-crossing partition.
2020 Mathematics Subject Classification:
Primary: 05C20, 46L53, 60F05; Secondary: 06A07, 46L53, 60E07.
1. Introduction
Non-commutative probability is based on various notions of independence for non-commuting random variables. The non-commuting variables are represented as elements of some unital *-algebra $\mathcal{A}$ (often an algebra of operators on a Hilbert space), and the expectation is represented by a state (a positive unital linear functional) $\phi:\mathcal{A}\to\mathbb{C}$. There are several notions of independence in the non-commutative setting. The most famous and fruitful is freeness, defined by Voiculescu [26, 27] (the same condition appeared also in Avitzour [3]). Muraki [17, 18] invented monotonic independence and in Bożejko’s paper [6] the condition for boolean independence appeared (for which Speicher and Woroudi [24] developed richer theory). It was show by Schürmann, Speicher and Muraki that there are only five universal notions of independence: classical one, freeness, monotone and antimonotone and boolean. On the other hand, there are several mixtures of these notions, which do not enjoy the universality property, nevertheless they allow to develop theories in analogy with classical probability. In particular, Wysoczański developed theory of bm-independence, which is a mixture of boolean and monotonic ones [29], Młotkowski [15] studied mixture of classical and free independence (under the name of $\Lambda$ independence), which was then developed by Speicher and Wysoczański [25]. A mixture of boolean and free independence, called bf-independence, was introduced by Kula and Wysoczański [11, 12]. Recently, Arizmendi, Rogelio Mendoza, and Vazquez-Becerra [1], introduced the notion of BMT independence (through a directed graph) as mixtures of boolean, monotone and tensor independences, and provided the corresponding central and Poisson-Type Limit Theorems.
Our work focuses on mixtures of free, boolean, and monotone independence that are described by directed graphs, as in [10, §3.2, 5.5], which extends both bf and bm independence of [11] and [29]. We aim to generalize the existing limit theorems in several ways:
  • We generalize the digraphs: We consider an arbitrary sequence of digraphs $G_{n}$ with number of vertices tending to infinity, requiring only the convergence of the normalized number of homomorphisms from trees into $G$. Examples of such graphs include both the discretizations of cones in [20, 19, 21] and the iterated compositions of a fixed graph from [10]. We also describe several new examples in §5.
  • We generalize the measures: We show that the $G_{n}$-free convolutions $\boxplus_{G_{n}}(\mu_{n})$ converge for any sequence of measures $\mu_{n}$ such that the boolean convolution powers $\mu_{n}^{\uplus|V_{n}|}$ converge. Thus, in particular, we obtain a limit theorem for each classical domain of attraction, in the spirit of Bercovici-Pata [4]. We deduce the general limit theorem from the compactly supported case using tools of [4, 9].
To state the results more precisely, first recall that a digraph is a pair $(V,E)$ where $V$ is the vertex set and the directed edge set $E\subseteq V\times V$ does not intersect the diagonal; in other words, $E$ viewed as a relation on $V$ is irreflexive. We write $v\rightsquigarrow w$ and $w\leftsquigarrow v$ when $(v,w)\in E$. Several other types of combinatorial objects important to our paper can be viewed as subclasses of digraphs:
  1. (1)
    A graph is a digraph $(V,E)$ such that $(v,w)\in E$ if and only if $(w,v)\in E$, or equivalently, $E$ is a symmetric relation on $V$.
  2. (2)
    A partially ordered set or poset is a digraph satisfying
    $v\rightsquigarrow w\implies w\not\rightsquigarrow v$
    and
    $v_{1}\rightsquigarrow v_{2}\text{ and }v_{2}\rightsquigarrow v_{3}\implies v_{1}\rightsquigarrow v_{3};$
    in other words, $E$ is antisymmetric and transitive, or $E$ is a strict partial order on $V$.
  3. (3)
    A rooted tree is a digraph $G=(V,E)$ such that there is some vertex $v$, called the root vertex such that for each $w\in V$, there exists a unique directed path from $v$ to $w$.
  4. (4)
    More generally, a rooted forest is a digraph such that there is some $S\subseteq V$ such that for each $w\in V$, there exists a unique directed path that starts in $S$ and ends at $w$.
If $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are digraphs, then a digraph homomorphism $\phi:G_{1}\to G_{2}$ is a map $V_{1}\to V_{2}$ such that if $v\rightsquigarrow w$ in $G_{1}$, then $\phi(v)\rightsquigarrow\phi(w)$ in $G_{2}$. We denote the set of homomorphisms by $\operatorname{Hom}(G_{1},G_{2})$.
For each finite digraph $G$, there is an associated convolution operation $\boxplus_{G}:\mathcal{P}(\mathbb{R})^{V}\to\mathcal{P}(\mathbb{R})$, where $\mathcal{P}(\mathbb{R})$ denotes the space of Borel probability measures on $\mathbb{R}$. If the input measures $\mu_{v}$ have compact support, they can be viewed as spectral distributions of bounded operators $X_{n}$, and then the convolution operation is defined by creating $G$-independent copies of $X_{n}$ though the explicit Hilbert space construction described in [10], and in §3.1 below. The convolution can also be described by calculating its moments; see Theorem 3.11 and Lemma 4.1 below. For the case of measures $\mu_{v}$ with unbounded support, the convolution can be described complex-analytically, or obtained by writing the measures $\mu_{v}$ as weak-* limits of compactly supported probability measures (see §4.2).
Our main result is stated as follows.
Theorem 1.1.
Let $G_{n}=(V_{n},E_{n})$ be a sequence of finite digraphs such that $\lim_{n\to\infty}|V_{n}|=\infty$. Suppose that for every finite rooted tree $G^{\prime}=(V^{\prime},E^{\prime})$, the limit
$\beta_{G^{\prime}}:=\lim_{n\to\infty}\frac{1}{|V_{n}|^{|V^{\prime}|}}|\operatorname{Hom}(G^{\prime},G_{n})|\text{ exists.}$.
Let $\mu\in\mathcal{P}(\mathbb{R})$ and let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence of probability measures on $\mathbb{R}$ such that $\lim_{n\to\infty}\mu_{n}^{\uplus|V_{n}|}=\mu$. Then
$\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})\text{ exists,}$,
where $\mu^{\uplus t}=\mu_{t}$ denotes the $t$-trasformation of $\mu$, introduced and studied in [7, 8] and shown there to be the boolean convolution power for $t\in(0,\infty)$. Furthermore, $\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})$ depends only upon $\mu$ and the coefficients $\beta_{G^{\prime}}$ for finite rooted trees $G^{\prime}$.
The proof of this theorem (see §4) proceeds in two stages. We first show the result for compactly supported measures using moments computations. Then we extend it to arbitrary measures using the results of [9].
Next, we turn to applications of the main theorem, exhibiting several classes of examples where the limits exist. Many of these results are obtained by viewing the digraphs $G_{n}$ as discretizations of some continuum object, which is a measurable digraph $(\Omega,\rho,S)$, that is, a complete probability measure space $(\Omega,\rho)$ representing some set of vertices, and a measurable $\mathcal{E}\subseteq\Omega\times\Omega$ representing the set of directed edges. As a concrete example, the reader may think of $\Omega$ being $[0,1]^{d}$ and $\mathcal{E}$ as some subset of $[0,1]^{2d}$ defined by inequalities. One may approximate $(\Omega,S)$ by discrete digraphs $G_{n}$ by partitioning $\Omega$ into measurable subsets $(A_{n,v})_{v\in V_{n}}$ of measure $1/|V_{n}|$ and choosing a subset $E_{n}\subseteq V_{n}\times V_{n}$ such that $\tilde{E}_{n}=\bigcup_{(v,w)\in E_{n}}A_{v}\times A_{w}$ converges to $\mathcal{E}$ in measure as $n\to\infty$. In this case, the $G_{n}$’s will satisfy the hypotheses of Theorem 1.1. More precisely, we have the following result (see §5).
Proposition 1.2.
Let $(\Omega,\rho)$ be a complete probability measure space and let $\mathcal{E}\subseteq\Omega\times\Omega$ be measurable. For each $n\in\mathbb{N}$, let $G_{n}=(V_{n},E_{n})$ be a finite digraph. Let $(A_{n,v})_{v\in V_{n}}$ be a measurable partition of $\Omega$ into sets of measure $1/|V_{n}|$, and let ~$\tilde{\mathcal{E}}_{n}=\bigcup_{(v,w)\in E_{n}}A_{n,v}\times A_{n,w}$. Suppose that $\rho(\tilde{\mathcal{E}}_{n}\Delta\mathcal{E})\to 0$. Then for every digraph $G^{\prime}=(V^{\prime},E^{\prime})$, we have
$\lim_{n\to\infty}\frac{\operatorname{Hom}(G^{\prime},G_{n})}{|V_{n}|^{|V^{\prime}|}}=\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,S))).$
In particular, if $\mu\in\mathcal{P}(\mathbb{R})$ and $(\mu_{n})_{n\in\mathbb{N}}$ is a sequence of probability measures on $\mathbb{R}$ such that $\lim_{n\to\infty}\mu_{n}^{\uplus|V_{n}|}=\mu$, then $\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})$ exists.
Using Proposition 1.2 and similar techniques, we show how Theorem 1.1 applies in several families of examples:
  1. (1)
    BM-independences described by symmetric cones as in [12, 19].
  2. (2)
    Iterated compositions of the same digraph in the sense of [10].
  3. (3)
    Regular graphs and more generally multi-regular graphs.
  4. (4)
    Sparse graphs.
Given that the coefficients $\beta_{G^{\prime}}$ can often be described as measures of the set of homomorphisms into some measurable digraph, it is natural to model the limiting measures using operators on a continuum analog of the $G$-free product Hilbert space. This leads to the construction of the Fock space associated to $(\Omega,S)$, which generalizes the BF and BM Fock spaces in [11, 20] and overlaps with [10] (and of course encompasses Fock spaces in the free [26], boolean, and monotone case [14, 16]).
This paper is organized as follows. In section 2, we present elementary background on non-commutative probability spaces, digraphs, and non-crossing partitions. In Section 3, we give a self-contained explanation of the Hilbert space construction and moment formulas for $G$-free independence. In §4, we prove Theorem 1.1. In §5, we prove Proposition 1.2 and study several families of examples. In §6, we describe the Fock space construction associated to continuum digraphs.
Acknowledgements
We thank the organizers of the International Workshop on Operator Theory and its Applications (Lisbon, Portugal 2019) and of the 19th Workshop: Noncommutative probability, noncommutative harmonic analysis and related topics, with applications, (Bedlewo, Poland 2022) which facilitated collaboration among the authors. We also thank the Mathematische Forschungsinstitut Oberwolfach and the University of Wrocław for funding DJ’s visit to Wrocław in May 2024. DJ thanks Weihua Liu, Ethan Davis, and Zhichao Wang for past collaboration on related projects.
2. Preliminaries
In this section, we review some basic knowledge and definitions for our study.
2.1. Non-commutative probability spaces
Here a non-commutative probability space refers to a unital $\mathrm{C}^{*}$-algebra $\mathcal{A}$ with a state $\phi:\mathcal{A}\to\mathbb{C}$ such that $\phi(bac)=0$ for all $b,c\in\mathcal{A}$ implies that $a=0$.
An important example is when $\mathcal{A}=B(\mathcal{H})$ for some Hilbert space $\mathcal{H}$, and $\phi:\mathcal{A}\to\mathbb{C}$ is given by $\phi(T)=\langle\xi,T\xi\rangle$ for some unit vector $\xi\in\mathcal{H}$. In fact, for every non-commutative probability space $(\mathcal{A},\phi)$, there exists a Hilbert space $\mathcal{H}$, unit vector $\xi$, and injective *-homomorphism $\iota:\mathcal{A}\to B(\mathcal{H})$ such that $\phi(a)=\langle\xi,\iota(a)\xi\rangle$ for all $a\in\mathcal{A}$.
2.2. Non-crossing partitions
Non-crossing partitions are a combinatorial tool that has been used to describe moments in non-commutative probability since the work of Speicher [22, 23]. We recall the relevant definitions and facts here.
We use the notation $[k]=\{1,\dots,k\}$. A partition of $[k]$ is a collection $\pi$ of subsets of $[k]$ called blocks such that $[k]=\bigsqcup_{B\in\pi}B$.
Definition 2.1 (Non-crossing).
For a partition $\pi$ of $[k]$, a crossing is a sequence of indices $i<i^{\prime}<j<j^{\prime}$ such that $i$ and $j$ are in some block $B$ and $i^{\prime}$ and $j^{\prime}$ are in some block $B^{\prime}\neq B$. We say that $\pi$ is non-crossing if it has no crossings. We denote by $\mathcal{NC}_{k}$ the set of non-crossing partitions of $[k]$.
Remark 2.2.
Visually, a partition is non-crossing if, after arranging points labeled $1$, …, $k$ on the circumference of a circle, it is possible to connect all the points in the same partition by curves passing through the disk such that the curves associated to points in two different blocks never cross each other.
Definition 2.3.
If $\pi$ is a partition of $[k]$ and $B^{\prime},B\in\pi$, we say that $B^{\prime}$ is nested inside $B$ if there exist $i,j\in B$ such that $i<j$ and $B^{\prime}\subseteq\{i+1,\dots,j-1\}\subseteq[k]\setminus B$; in other words, there are no intervening indices of $B$ between $i$ and $j$, and $B^{\prime}$ lies entirely between $i$ and $j$. In this case, we write $B\prec B^{\prime}$.
Definition 2.4.
If $\pi$ is a partition of $[k]$ and $B,B^{\prime}\in\pi$, we say that $B$ and $B^{\prime}$ are separated if there exists $j\in[k]$ such that either $B\subseteq\{1,\dots,j\}$ and $B^{\prime}\subseteq\{j+1,\dots,k\}$ or $B^{\prime}\subseteq\{1,\dots,j\}$ and $B\subseteq\{j+1,\dots,k\}$.
Lemma 2.5.
Let $\pi$ be a non-crossing partition and $B$, $B^{\prime}$ are distinct blocks of $\pi$. Then either $B^{\prime}$ is nested inside $B$, $B$ is nested inside $B^{\prime}$, or $B$ and $B^{\prime}$ are separated, and these cases are mutually exclusive.
Proof.
Suppose that there exists $i,j\in B$ and $i^{\prime}\in B^{\prime}$ such that $i<i^{\prime}<j$. Without loss of generality, assume that $i$ is the largest index in $B$ to the left of $i^{\prime}$, and $j$ is the smallest index in $B$ to the right of $i^{\prime}$. Then $\{i+1,\dots,j-1\}\subseteq[k]\setminus B$. If $B^{\prime}$ had some element $j^{\prime}$ that was not contained in $\{i+1,\dots,j-1\}$, then $i<i^{\prime}<j<j^{\prime}$ would be a crossing. Hence, $B^{\prime}\subseteq\{i+1,\dots,j-1\}$, so $B^{\prime}$ is nested inside $B$.
Similarly, if there exists $i^{\prime},j^{\prime}\in B^{\prime}$ and $j\in B$ such that $i^{\prime}<j<j^{\prime}$, then $B$ is nested inside $B^{\prime}$.
If neither of the two cases above holds, then either all the indices of $B$ are less than those of $B^{\prime}$, or vice versa, hence $B$ and $B^{\prime}$ are separated. It is a straightforward exercise that $B\prec B^{\prime}$, $B^{\prime}\prec B$, and $B$ and $B^{\prime}$ separated are mutually exclusive cases. ∎
Corollary 2.6.
For $\pi\in\mathcal{NC}_{k}$, the nesting relation $\prec$ is a strict partial order on $\pi$.
Proof.
The previous lemma shows that $B\prec B^{\prime}$ and $B^{\prime}\prec B$ are mutually exclusive. It is immediate that $B\prec B^{\prime}$ implies $B\neq B^{\prime}$, and straightforward to check that $\prec$ is transitive. ∎
Thus, $(\pi,\prec)$ is a poset. Recall that for a poset, the covering relation is the relation $R$ given by $xRy$ if $x<y$ and there is no $z$ with $x<z<y$. In this case, we call $x$ a predecessor of $y$.
Definition 2.7.
For $\pi\in\mathcal{NC}_{k}$, let $\mathrm{F}(\pi)$ be the digraph with vertex set $\pi$ and edges given by the covering relation of $(\pi,\prec)$. That is, $B\rightsquigarrow B^{\prime}$ if $B\prec B^{\prime}$ and there is no $B^{\prime\prime}$ with $B\prec B^{\prime\prime}\prec B^{\prime}$.
Lemma 2.8.
Let $\pi\in\mathcal{NC}_{k}$ and $B\in\pi$. Then either $B$ is minimal with respect to $\prec$, or there is a unique $B^{\prime}$ such that $B^{\prime}\rightsquigarrow B$ in $\mathrm{F}(\pi)$. In particular, $\mathrm{F}(\pi)$ is a rooted forest (viewed as a digraph with edges oriented away from the root of each component).
Proof.
Suppose that $B$ is not minimal. Then $\{B^{\prime}:B^{\prime}\prec B\}$ is a finite poset and hence has a maximal element, so there exists some $B^{\prime}$ with $B^{\prime}\sim B$ in $\mathrm{F}(\pi)$. To show that this $B^{\prime}$ is unique, consider some other $B^{\prime\prime}$ with $B^{\prime\prime}\prec B$. By Lemma 2.5, either $B^{\prime}\prec B^{\prime\prime}$ or $B^{\prime\prime}\prec B^{\prime}$ or $B^{\prime}$ and $B^{\prime\prime}$ are separated. The case $B^{\prime}\prec B^{\prime\prime}\prec B$ cannot happen because we assumed that $B^{\prime}\sim B$. If $B^{\prime\prime}\prec B^{\prime}$, then we cannot have $B^{\prime}\sim B$. Finally, if $B^{\prime}$ and $B^{\prime\prime}$ are separated, then there exists a partition of $[k]$ into two intervals $I^{\prime}$ and $I^{\prime\prime}$ with $B^{\prime}\subseteq I^{\prime}$ and $B^{\prime\prime}\subseteq I^{\prime\prime}$. Since $B^{\prime}\prec B$ and $I^{\prime}$ is an interval, we have $B\subseteq I^{\prime}$. Hence, $B$ and $B^{\prime\prime}$ are separated, which contradicts $B^{\prime\prime}\prec B$, so the case where $B^{\prime}$ and $B^{\prime\prime}$ are separated also cannot happen. This completes the proof of the first claim.
To show that $\mathrm{F}(\pi)$ is a forest, one uses the first claim to construct a backward path from any given $B$ to some $B^{\prime}$ which is minimal in $\prec$ (that is, a directed path from $B^{\prime}$ to $B$) and check that this path is unique. ∎
Notation 2.9.
For $\pi\in\mathcal{NC}_{k}$, we call $\mathrm{F}(\pi)$ the nesting forest of $\pi$.
Notation 2.10.
We denote by $\operatorname{depth}(B)$ the depth of a block $B$ in the forest $\mathrm{F}(\pi)$. If $B$ is maximal, then $\operatorname{depth}(B)=1$. If $\operatorname{depth}(B)>1$, then we denote by $\operatorname{pred}(B)$ the predessor of $B$ (which is unique because $\mathrm{F}(\pi)$ is a forest).
3. $G$-free independence
Often a non-commutative probability paper defines independence of *-subalgebras $\mathcal{A}_{1}$, …, $\mathcal{A}_{n}$ first through a condition on moments and then uses a Hilbert space model to show that for any given non-commutative probability spaces $(\mathcal{A}_{j},\phi_{j})$, there exists some $(\mathcal{A},\phi)$ containing independent copies of $(\mathcal{A}_{1},\phi_{1})$, …, $(\mathcal{A}_{n},\phi_{n})$. However, since the moment formula for general digraphs (and even for mixtures of free and boolean independences [11]) is complicated, we will begin with the Hilbert space model, explain how the moment formula arises naturally from the Hilbert space structure, and use this for the definition of independence.
The digraph construction described here generalizes the BM-product Hilbert space in [21, §3.1]. Moreover, it is a special case of the more general tree construction of [10], corresponding to the case when the tree arises as the set of paths on the digraph. However, we want to present a self-contained explanation of the digraph case by itself since it admits a simpler notation and intuition.
3.1. Digraph products of pointed Hilbert spaces
Notation 3.1.
By a pointed Hilbert space, we mean a pair $(\mathcal{H},\xi)$ where $\mathcal{H}$ is a Hilbert space and $\xi\in\mathcal{H}$ is a unit vector. If $(\mathcal{H},\xi)$ is a pointed Hilbert space, we denote by $\mathcal{H}^{\circ}$ the orthogonal complement of $\mathbb{C}\xi$ in $\mathcal{H}$.
Notation 3.2.
Let $G=(V,E)$ be a digraph, and let
$E_{m}=\{(v_{0},\dots,v_{m}):v_{0}\rightsquigarrow v_{1}\rightsquigarrow v_{2}\rightsquigarrow\dots\rightsquigarrow v_{m}\}$
be the set of (directed) paths of length $m$. Note that $E_{0}=V$ and $E_{1}=E$. We also write
$E_{m}^{\dagger}=\{(v_{0},\dots,v_{m}):v_{0}\leftsquigarrow v_{1}\leftsquigarrow v_{2}\leftsquigarrow\dots\leftsquigarrow v_{m}\}$
for the set of reverse paths.
Definition 3.3 ($G$-free product of pointed Hilbert spaces).
Let $G=(V,E)$ be a digraph, and let $(\mathcal{H}_{v},\xi_{v})_{v\in V}$ be a collection of pointed Hilbert spaces indexed by $V$. We define 🟊$\bigstar_{G}[(\mathcal{H}_{v},\xi_{v})_{v\in V}]$ as the pointed Hilbert space $(\mathcal{H},\xi)$ given by
(3.1) $\mathcal{H}=\mathbb{C}\xi\oplus\bigoplus_{m\geq 0}\bigoplus_{(v_{0},\dots,v_{m})\in E_{m}^{\dagger}}\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}^{\circ}.$
Here we can think of $\mathcal{H}_{v}$ as sitting inside $\mathcal{H}$ by identifying $\xi_{v}$ with $\xi$.
Definition 3.4.
Continuing with the notation of the previous definition, we define for each $v\in V$ a *-homomorphism $\iota_{v}:B(\mathcal{H}_{v})\to B(\mathcal{H})$ as follows. Let
$\mathcal{H}_{\rightsquigarrow v}=\mathbb{C}\xi\oplus\bigoplus_{m\geq 0}\bigoplus_{\begin{subarray}{c}(v_{0},\dots,v_{m})\in E_{m}^{\dagger}v_{0}\rightsquigarrow v\end{subarray}}\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$
and
$\mathcal{H}_{\perp v}=\bigoplus_{m\geq 0}\bigoplus_{\begin{subarray}{c}(v_{0},\dots,v_{m})\in E_{m}^{\dagger}\\ v_{0}\neq v\\ v_{0}\text{ not }\rightsquigarrow v\end{subarray}}\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}.$
By distributing tensor products over direct sums, we have an unitary isomorphism
(3.2) $u_{v}:\mathcal{H}\to[(\mathbb{C}\oplus\mathcal{H}_{v}^{\circ})\otimes\mathcal{H}_{\rightsquigarrow v}]\oplus\mathcal{H}_{\perp v}\to[\mathcal{H}_{v}\otimes\mathcal{H}_{\rightsquigarrow v}]\oplus\mathcal{H}_{\perp v},$
where the first term $\mathbb{C}\otimes\mathcal{H}_{\rightsquigarrow v}=\mathcal{H}_{\rightsquigarrow v}$ corresponds to reverse paths that start with a vertex $v^{\prime}\rightsquigarrow v$, the second term $\mathcal{H}_{v}^{\circ}\otimes\mathcal{H}_{\rightsquigarrow v}$ corresponds to reverse paths that start with $v$, and the third term $\mathcal{H}_{\perp v}$ corresponds to all other reverse paths. Then we define
$\iota_{v}(a)=u_{v}\left([a\otimes\operatorname{id}_{\mathcal{H}_{\rightsquigarrow v}}]\oplus 0_{\mathcal{H}_{\perp V}}\right)u_{v}^{*}\text{ for }a\in B(\mathcal{H}_{v}).$
The next lemma shows that $\iota_{v}$ is expectation-preserving.
Lemma 3.5.
With the setup and notation of the previous two definitions,
$\langle\xi,\iota_{v}(a)\xi\rangle=\langle\xi_{v},a\xi_{v}\rangle\text{ for all}a\in B(\mathcal{H}_{v}).$
Proof.
The subspace $\mathbb{C}\xi\oplus\mathcal{H}_{v}^{\circ}\cong\mathcal{H}_{v}$ of $\mathcal{H}$ is contained corresponds to the term $\mathcal{H}_{v}\otimes\mathbb{C}\subseteq\mathcal{H}_{v}\otimes\mathcal{H}_{\rightsquigarrow v}$ in (3.2). From this we see that $\mathbb{C}\xi\oplus\mathcal{H}_{v}^{\circ}$ is an invariant subspace of $\iota_{v}(a)$ on which $\iota_{v}(a)$ acts in the same way as $a$ acts on $\mathcal{H}_{v}=\mathbb{C}\xi_{v}\oplus\mathcal{H}_{v}^{\circ}$. The conclusion of the lemma is immediate from this. ∎
3.2. Computation of joint moments
Let $G=(V,E)$ be a digraph and let $(\mathcal{H}_{v},\xi_{v})_{v\in V}$ be a family of pointed Hilbert spaces indexed by $V$. Let $(\mathcal{H},\xi)$ be the $G$-free product Hilbert space (Definition 3.3) and let $\iota_{v}:B(\mathcal{H}_{v})\to B(\mathcal{H})$ be the *-homomorphisms described in Definition 3.4. Our goal is to compute
$\langle\xi,\iota_{\ell(1)}(a_{1})\dots\iota_{\ell(k)}(a_{k})\xi\rangle,$
where $\ell:[k]\to V$ is a function and $a_{j}\in B(\mathcal{H}_{\ell(j)})$. The resulting moment formula will be the basis for the definition of $G$-free independence.
In order to state this formula, we will use non-crossing partitions as well as boolean cumulants, which are defined as follows.
Definition 3.6.
A partition $\pi$ of $[k]$ is called an interval partition if every block $B\in\pi$ has the form $B=\{i,\dots,j\}$ for some $1\leq i\leq j\leq k$. We denote by $I_{k}$ the set of interval partitions of $[k]$.
Definition 3.7.
Let $(\mathcal{A},\phi)$ be a non-commutative probability space. We define $K_{\operatorname{bool},k}:\mathcal{A}^{k}\to\mathbb{C}$ by
$K_{\operatorname{bool},k}[a_{1},\dots,a_{k}]=\sum_{\pi\in I_{k}}(-1)^{|\pi|-1}\prod_{B\in\pi}\phi\left(\prod_{j\in B}a_{j}\right),$
where $\prod_{j\in B}a_{j}$ denotes the product of the $a_{j}$ for $j\in B$ written in order from left to right.
Lemma 3.8.
Let $(\mathcal{A},\phi)$ be a non-commutative probability space and $(\mathcal{H},\xi)$ a pointed Hilbert space with $\mathcal{A}\subseteq B(\mathcal{H})$ and $\phi(a)=\langle\xi,a\xi\rangle$. Let $P\in B(\mathcal{H})$ be the rank-one projection onto $\xi$ and $Q=1-P$. Then
$K_{\operatorname{bool},k}[a_{1},\dots,a_{k}]=\langle\xi,a_{1}Qa_{2}\dots Qa_{k}\xi\rangle.$
Proof.
Write
$\langle\xi,a_{1}Qa_{2}\dots Qa_{k}\xi\rangle$ $=\langle\xi,a_{1}(1-P)a_{2}\dots(1-P)a_{k}\xi\rangle$
$=\sum_{T_{1},\dots,T_{k-1}\in\{1,-P\}}\langle\xi,a_{1}T_{1}a_{2}\dots T_{k-1}a_{k}\xi\rangle.$
Consider the map from the set $I_{k}$ of interval partitions to the sequences $T_{1}$, …, $T_{k}$ from $\{1,-P\}$ that sends $\pi\in I_{k}$ to the sequence $T_{1}$, …, $T_{k}$ with $T_{j}=1$ if $j$ and $j+1$ are in the same block of $\pi$ and $T_{j}=-P$ if $j$ and $j+1$ are in different blocks of $\pi$. It is straightforward to check that this is a bijection. Moreover, if $T_{1}$, …, $T_{k}$ is the sequence associated to $\pi$, and if $j_{1},\dots,j_{|\pi|-1}$ are the indices where $j$ and $j+1$ are in distinct blocks, then we have
$\langle\xi,a_{1}T_{1}a_{2}\dots T_{k-1}a_{k}\xi\rangle$ $=(-1)^{|\pi|-1}\left\langle\xi,a_{1}\dots a_{j_{1}}Pa_{j_{1}+1}\dots a_{j_{2}}\dots Pa_{j_{|\pi|-1}}\dots a_{k}\xi\right\rangle$
$=(-1)^{|\pi|-1}\langle\xi,a_{1}\dots a_{j_{1}}\xi\rangle\dots\langle\xi,a_{j_{|\pi|-1}}\dots a_{k}\xi\rangle$
$=(-1)^{|\pi|-1}\prod_{B\in\pi}\phi\left(\prod_{j\in B}a_{j}\right).$
This is precisely the definition of $K_{\operatorname{bool},k}[a_{1},\dots,a_{k}]$, so the proof is complete. ∎
Definition 3.9.
Let $G=(V,E)$ be a digraph. By a $V$-labelling of $[k]$, we mean a function $\ell:[k]\to V$. For every such labelling, we denote by $\mathcal{NC}_{k}(\ell)$ the set of $\pi\in\mathcal{NC}_{k}$ such that $\ell$ is constant on each block of $\pi$; we say that $\pi$ and $\ell$ are compatible if $\pi\in\mathcal{N}C_{k}(\ell)$.
Definition 3.10.
Let $G=(V,E)$ be a digraph and $\ell:[k]\to V$. Let $\pi\in\mathcal{NC}_{k}(\ell)$. Let ~$\tilde{\ell}:\pi\to V$ be the map given by $\ell(B)=\{\tilde{\ell}(B)\}$. We define $\mathcal{NC}_{k}(\ell,G)$ as the set of $\pi\in\mathcal{NC}_{k}(\ell)$ such that $\tilde{\ell}$ defines a digraph homomorphism $\mathrm{F}(\pi)\to G$. In this case, we say that $\pi$, $\ell$, and $G$ are compatible.
Theorem 3.11.
Let $G=(V,E)$ be a digraph. Let $(\mathcal{H}_{v},\xi_{v})_{v\in V}$ be a collection of pointed Hilbert spaces indexed by $v\in V$, let $(\mathcal{H},\xi)$ be the $G$-free product as in Definition 3.3, and let $\iota_{v}:B(\mathcal{H}_{v})\to B(\mathcal{H})$ for $v\in V$ be the *-homomorphisms given in Definition 3.4. Let
$\phi_{v}:$ $B(\mathcal{H}_{v})\to\mathbb{C}:a\mapsto\langle\xi_{v},a\xi_{v}\rangle$
ϕ::italic-$\phi:$ $B(\mathcal{H})\to\mathbb{C}:a\mapsto\langle\xi,a\xi\rangle,$
so that $(B(\mathcal{H}_{v}),\phi_{v})$ and $(B(\mathcal{H}),\phi)$ are non-commutative probability spaces and $\iota_{v}:B(\mathcal{H}_{v})\to B(\mathcal{H})$ is expectation-preserving by Lemma 3.5.
Let $k\in\mathbb{N}$, let $\ell:[k]\to V$ be a labelling, and let $a_{j}\in B(\mathcal{H}_{\ell(j)})$ for $j=1$, …, $k$. Then
$\langle\xi,\iota_{\ell(1)}(a_{1})\dots\iota_{\ell(k)}(a_{k})\xi\rangle=\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[a_{j}:j\in B].$
where for each block $B$, $K_{\operatorname{bool},|B|}$ denotes the $|B|$th boolean cumulant associated to $(B(\mathcal{H}_{\ell(B)}),\phi_{\ell(B)})$, and the arguments $a_{j}:j\in B$ are written in increasing order of their indices from left to right.
Remark 3.12.
For each block $B$ of a partition $\pi$ in the above formula, since the map $\iota_{\ell(B)}$ is expectation-preserving, we could equivalently write
$\langle\xi,\iota_{\ell(1)}(a_{1})\dots\iota_{\ell(k)}(a_{k})\xi\rangle=\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[\iota_{\ell(j)}(a_{j}):j\in B].$
Proof of Theorem 3.11.
As a notational convenience, let us reindex the operators $\iota_{\ell(1)}(a_{1})$, …, $\iota_{\ell(k)}(a_{k})$ in reverse order, so that $\iota_{\ell(1)}(a_{1})$ is the right-most operator, i.e. it is applied to $\xi$ first. Thus, we want to prove that
(3.3) $\langle\xi,\iota_{\ell(k)}(a_{k})\dots\iota_{\ell(1)}(a_{1})\xi\rangle=\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[a_{j}:j\in B],$
with the indices $a_{j}:j\in B$ now in decreasing order for each block.
Let $P_{v}\in B(\mathcal{H}_{v})$ be the rank-one projection onto $\mathbb{C}\xi_{v}$, and let $Q_{v}=1-P_{v}$. Let
$a_{j}^{(0,0)}$ $=P_{\ell(j)}a_{j}P_{\ell(j)}$
$a_{j}^{(0,1)}$ $=P_{\ell(j)}a_{j}Q_{\ell(j)}$
$a_{j}^{(1,0)}$ $=Q_{\ell(j)}a_{j}P_{\ell(j)}$
$a_{j}^{(1,1)}$ $=Q_{\ell(j)}a_{j}Q_{\ell(j)}.$
Then $a_{j}=a_{j}^{(0,0)}+a_{j}^{(0,1)}+a_{j}^{(1,0)}+a_{j}^{(1,1)}$. We may thus write
(3.4) $\langle\xi,\iota_{\ell(k)}(a_{k})\dots\iota_{\ell(1)}(a_{1})\xi\rangle=\sum_{\delta_{1},\epsilon_{1},\dots,\delta_{k},\epsilon_{k}\in\{0,1\}}\langle\xi,\iota_{\ell(k)}(a_{k}^{(\delta_{k},\epsilon_{k})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi\rangle.$
Our goal is to show that certain of the terms in the sum vanish, while the others correspond to non-crossing partitions and evaluate to the product of boolean cumulants in the asserted formula. Note that $a_{j}^{(\delta,\epsilon)}$ annihilates $\mathcal{H}_{\ell(j)}^{\circ}$ when $\epsilon=0$ and annihilates $\mathbb{C}\xi_{\ell(j)}$ when $\epsilon=1$, and its image is contained in $\mathbb{C}\xi_{\ell(j)}$ when $\delta=0$ and $\mathcal{H}_{\ell(j)}^{\circ}$ when $\delta=1$.
Examining the definition of the maps $\iota_{v}$ in Definition 3.4, we conclude the following.
Observation 3.13.
  • $\iota_{\ell(j)}(a^{(0,0)})$ maps $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ into itself if $v_{0}\sim\ell(j)$, and vanishes on $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ otherwise.
  • $\iota_{\ell(j)}(a^{(1,0)})$ maps $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ into $\mathcal{H}_{\ell(j)}^{\circ}\otimes\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ if $v_{0}\sim\ell(j)$, and vanishes on $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ otherwise.
  • $\iota_{\ell(j)}(a^{(0,1)})$ maps $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ into itself if $v_{0}\sim\ell(j)$, and vanishes on $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ otherwise.
  • $\iota_{\ell(j)}(a^{(1,1)})$ maps $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ into itself if $v_{0}=\ell(j)$, and vanishes on $\mathcal{H}_{v_{0}}^{\circ}\otimes\dots\otimes\mathcal{H}_{v_{m}}$ otherwise.
With this information in mind, we can then consider the effect of applying several operators $\iota_{\ell(j)}(a^{(\delta_{j},\epsilon_{j})})$ consecutively to the state vector $\xi$, and thus determine which direct summand of the Hilbert space $\mathcal{H}$ contains the vector
$\iota_{\ell(j)}(a_{j}^{(\delta_{1},\epsilon_{1})})\dots\iota_{\ell(k)}(a_{k}^{(\delta_{k},\epsilon_{k})})\xi$
for each $j\leq k$. First, to keep track of the number of tensorands, we introduce a height function $h$ associated to the sequence of indices $(\delta_{j},\epsilon_{j})$. Let
$h(m)=\sum_{j=1}^{m}(\delta_{i}-\epsilon_{i}).$
Note that $h(0)=0$, and $h(j+1)-h(j)\in\{-1,0,1\}$. By inductive application of the observations above, one can show that $\iota_{\ell(j)}(a_{j})\dots\iota_{\ell(1)}(a_{1})\xi$ is contained in one of the $h(j)$-fold tensor products among the direct summands in the definition of $\mathcal{H}$, provided that $h(i)\geq 0$ for $i\leq j$. If $h(j)$ is ever $-1$, then the first time that $h(j)=-1$, we are applying an ‘annihilation operator’ $\iota_{\ell(j)}(a_{j}^{(0,1)})$ to a multiple of the state vector $\xi$, which results in $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi=0$. Hence also, if $h(i)<0$ for any $i\leq j$, then $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi=0$. Furthermore, at the last step, for the inner product to be nonzero, $\iota_{\ell(k)}(a_{k}^{(\delta_{k},\epsilon_{k})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi$ must be in $\mathbb{C}\xi$, and hence $h(k)=0$.
Therefore, in the expansion 3.4, only the summands which have a nonnegative height function $h$ with $h(k)=0$ will remain. We want to express these in terms of non-crossing partitions. Thus, we recall the following fact. This is a generalization of the well-known bijection between non-crossing pair partitions and Dyck paths. We will not give the proof here in detail, since a similar argument is given in [10, Lemma 4.24]. However, note here that we are picturing the indices $1$, …, $k$ as running from right to left.
Lemma 3.14.
There is a bijection between
  1. (1)
    sequences $(\delta_{1},\epsilon_{1})$, …, $(\delta_{k},\epsilon_{k})$ whose height function $h$ is nonnegative and satisfies $h(k)=0$, and
  2. (2)
    non-crossing partitions $\pi\in\mathcal{NC}_{k}$,
described by the following relationship:
  • $\{j\}$ is a singleton in $\pi$ if and only if $(\delta_{j},\epsilon_{j})=(0,0)$.
  • $\{j\}$ is the upper (left) endpoint of a non-singleton block in $\pi$ if and only if $(\delta_{j},\epsilon_{j})=(0,1)$.
  • $\{j\}$ is the lower (right) endpoint of a non-singleton block in $\pi$ if and only if $(\delta_{j},\epsilon_{j})=(1,0)$.
  • $\{j\}$ is in a non-singleton block and not an endpoint of the block if and only if $(\delta_{j},\epsilon_{j})=(1,1)$.
Now given a non-crossing partition $\pi$, we need to evaluate the corresponding term $\langle\xi,\iota_{\ell(k)}(a_{k}^{(\delta_{k},\epsilon_{k})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi\rangle$. In particular, we must show it is zero unless $\pi$, $\ell$, and $G$ are compatible.
We aim to evaluate $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi$ by induction on $j$. To this end, we introduce more notation. Let $\pi_{j}$ be the restriction of $\pi$ to $[j]$, which is a non-crossing partition. Each block of $\pi_{j}$ is thus $B\cap[j]$ for some block $B$ in $\pi$. A block of $\pi_{j}$, say $B\cap[j]$, is called finished if $B\cap[j]=B$ and unfinished otherwise. Let $\mathrm{F}_{j}$ be the forest where there is an edge from $B\cap[j]$ to $B^{\prime}\cap[j]$ in $\mathrm{F}_{j}$ if and only if there is an edge from $B$ to $B^{\prime}$ in $\mathrm{F}(\pi)$, which is a subgraph of $\mathrm{F}(\pi)$.
Lemma 3.15.
If the labelling $\ell$ is constant on each block of $\pi_{j}$ and defines a homomorphism from $\mathrm{F}_{j}$ to $G$, then
(3.5) $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi\\ =\bigotimes_{r=m}^{1}\left[\left(\prod_{s\in B_{r}}Q_{\ell(B_{r})}a_{s}\right)\xi_{\ell(B_{r})}\right]\prod_{B\in\pi_{j}\text{ finished}}K_{\operatorname{bool},|B|}[a_{s}:s\in B].$
where $B_{1}$, …, $B_{m}$ are the unfinished blocks of $\pi_{j}$, ordered by $\min B_{1}<\dots<\min B_{m}$, and the terms in $\prod_{s\in B_{r}}Q_{\ell(B_{r})}a_{s}$ are multiplied from left to right in decreasing order of the index $s$. Here, as above in (3.3), the terms $a_{s}:s\in B$ in the boolean cumulant also run in decreasing order from left to right. In all other cases, $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi=0$.
Proof.
We proceed by induction. The base case $j=0$ is immediate; all the products are empty and so both sides evaluate to $\xi$. For the induction step, suppose the claim is true for $j$ and we will prove it for $j+1$. For simplicity, let us denote by (*) the condition that the labelling $\ell$ is constant on each block of $\pi_{j}$ and defines a homomorphism from $\mathrm{F}_{j}$ to $G$.
If (*) fails for $j$, then it also fails for $j+1$. By induction hypothesis, $\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi=0$, and hence also $\iota_{\ell(j+1)}(a_{j+1}^{(\delta_{j+1},\epsilon_{j+1})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi=0$. Thus, the claim holds for $j+1$.
Now suppose that (*) holds for $j$. Note that
$\zeta_{j}:=\iota_{\ell(j)}(a_{j}^{(\delta_{j},\epsilon_{j})})\dots\iota_{\ell(1)}(a_{1}^{(\delta_{1},\epsilon_{1})})\xi\in\bigotimes_{r=m}^{1}\mathcal{H}_{\ell(B_{r})}^{\circ}.$
Using the same notation as (3.5) for the unfinished blocks, express $B_{1}$, …, $B_{m}$ as $[j]\cap B_{1}^{\prime}$, …, $[j]\cap B_{m}^{\prime}$ for blocks $B_{1}^{\prime}$, …, $B_{m}^{\prime}$ in $\pi$. Note that $B_{s+1}^{\prime}$ is nested inside $B_{s}^{\prime}$ because $\min B_{s+1}>\min B_{s}$ but $B_{s}^{\prime}$ contains an element greater than $\min B_{s+1}^{\prime}$ because $B_{s}$ is unfinished in $\pi_{j}$. Similar elementary reasoning with non-crossing conditions shows that there is no block strictly between $B_{s+1}$ and $B_{s}$ in the nesting order, so that $B_{s}^{\prime}\sim B_{s+1}^{\prime}$ in $\mathrm{F}(\pi)$, hence also $B_{s}\sim B_{s+1}$ in $\mathrm{F}_{j}$.
We consider cases based on $(\delta_{j+1},\epsilon_{j+1})$.
  1. (1)
    Suppose $(\delta_{j+1},\epsilon_{j+1})=(0,0)$, so that $a^{(0,0)}$ is a multiple of $P_{\ell(j+1)}$. Thus, $\{j+1\}$ is a singleton block in $\pi$ that is nested inside $B_{m}^{\prime}$. Moreover, $\{j+1\}$ is a finished block in $\pi_{j+1}$, and it is the only new vertex in $\mathrm{F}_{j+1}$ that was not in $\mathrm{F}_{j}$. Thus, $\ell$ defines a homomorphism $\mathrm{F}_{j+1}\to G$ if and only if $\ell(B_{m})\sim\ell(j+1)$. Therefore, if $\ell(B_{m})\not\sim\ell(j+1)$, then (*) fails for $j+1$ and $\iota_{\ell(j+1)}(a^{(0,0)})\zeta_{j}=0$ by Observation 3.13. On the other hand, if $\ell(B_{m})\sim\ell(j+1)$, then since $\zeta_{j}\in\bigotimes_{j=m}^{1}\mathcal{H}_{\ell(B_{j})}^{\circ}$, we obtain $\iota_{\ell(j+1)}(a_{j+1}^{(0,0)})\zeta_{j}=K_{\operatorname{bool},1}(a_{j+1})\zeta_{j}$; meanwhile, on the right-hand side of (3.5), a new term of $K_{\operatorname{bool},1}(a_{j+1})$ is added for the new finished block $\{j+1\}$ in $\pi_{j+1}$.
  2. (2)
    Suppose $(\delta_{j+1},\epsilon_{j+1})=(1,0)$. In this case $\{j+1\}$ is a singleton block in $\pi_{j+1}$ that is the right endpoint of a block in $\pi$. Similar to case (1), $\ell$ defines a homomorphism $\mathrm{F}_{j+1}\to V$ if and only if $\ell(B_{m})\sim\ell(j+1)$. Therefore, if $\ell(B_{m})\not\sim\ell(j+1)$, then (*) fails for $j+1$ and $\iota_{\ell(j+1)}(a^{(1,0)})\zeta_{j}=0$ by Observation 3.13. On the other hand, if $\ell(B_{m})\sim\ell(j+1)$, then since $\zeta_{j}\in\bigotimes_{j=m}^{1}\mathcal{H}_{\ell(B_{j})}^{\circ}$, we obtain $\iota_{\ell(j+1)}(a_{j+1}^{(1,0)})\zeta_{j}=a_{j+1}\xi_{\ell(j+1)}\otimes\zeta_{j}$; meanwhile, on the right-hand side of (3.5), a new term of $a_{j+1}\xi_{\ell(j+1)}$ is added in the tensor product expansion corresponding to the new unfinished block $\{j+1\}$ in $\pi_{j+1}$.
  3. (3)
    Suppose that $(\delta_{j+1},\epsilon_{j+1})=(1,1)$. In this case, $j+1$ is added to the most recent unfinished block $B_{m}$ in $\pi_{j}$. Thus, $\ell$ defines a homomorphism $\mathrm{F}_{j+1}\to V$ if and only if $\ell(j+1)=\ell(B_{m})$. If $\ell(j+1)\neq\ell(B_{m})$, then $\iota_{\ell(j+1)}(a_{j+1}^{(1,1)})\zeta_{j}=0$ by Observation 3.13. On the other hand, if $\ell(j+1)\neq\ell(B_{m})$, then
    $\iota_{\ell(j+1)}(a_{j+1}^{(1,1)})\bigotimes_{r=m}^{1}\left[\left(\prod_{s\in B_{m}}Q_{\ell(B_{m})}a_{s}\right)\xi_{\ell(B_{r})}\right]\\ =Q_{\ell(j+1)}a_{j+1}\left(\prod_{s\in B_{m}}Q_{\ell(B_{m})}a_{s}\right)\xi_{\ell(B_{m})}\bigotimes_{r=m-1}^{1}\left[\left(\prod_{s\in B_{r}}Q_{\ell(B_{r})}a_{s}\right)\xi_{\ell(B_{r})}\right].$
    This change is accounted for on the right-hand side of (3.5) by adding a new term corresponding to $j+1$ onto the product of $a_{s}$’s for the block $B_{m}$.
  4. (4)
    Suppose that $(\delta_{j+1},\epsilon_{j+1})=(1,1)$. Similar to case (3), $j+1$ is added to the most recent unfinished block $B_{m}$ in $\pi_{j}$, and this block is now finished in $\pi_{j+1}$. Thus, $\ell$ defines a homomorphism $\mathrm{F}_{j+1}\to V$ if and only if $\ell(j+1)=\ell(B_{m})$. If $\ell(j+1)\neq\ell(B_{m})$, then $\iota_{\ell(j+1)}(a_{j+1}^{(1,1)})\zeta_{j}=0$ by Observation 3.13. On the other hand, if $\ell(j+1)\neq\ell(B_{m})$, then
    $\iota_{\ell(j+1)}(a_{j+1}^{(0,1)})\bigotimes_{r=m}^{1}\left[\left(\prod_{s\in B_{m}}Q_{\ell(B_{m})}a_{s}\right)\xi_{\ell(B_{r})}\right]\\ =\left\langle\xi_{\ell(j+1)},a_{j+1}\left(\prod_{s\in B_{m}}Q_{\ell(B_{m})}a_{s}\right)\xi_{\ell(B_{m})}\right\rangle\bigotimes_{r=m-1}^{1}\left[\left(\prod_{s\in B_{r}}Q_{\ell(B_{r})}a_{s}\right)\xi_{\ell(B_{r})}\right]\\ =K_{\operatorname{bool},|B_{m}|+1}[a_{s}:s\in B_{m}\cup\{j+1\}]\bigotimes_{r=m-1}^{1}\left[\left(\prod_{s\in B_{r}}Q_{\ell(B_{r})}a_{s}\right)\xi_{\ell(B_{r})}\right].$
    This change is accounted for on the right-hand side of (3.5) by removing the block $B_{m}$ from the tensor product expansion for the unfinished blocks, and adding a new term for $B_{1}\cup\{j+1\}$ in the product expansion for the finished blocks.
In each case, the induction proceeds and completes the proof of the lemma. ∎
Now looking at the result of the lemma in the case where $j=k$, there are no unfinished blocks, and hence no tensor product terms. Thus, (3.5) reduces to $\prod_{B\in\pi}K_{\operatorname{bool},|B|}[a_{j}:j\in B]$. Thus, by Lemmas 3.14 and 3.15, the terms that survive in (3.4) correspond to partitions $\pi$ that are compatible with $\ell$ and $G$. Therefore, we obtain (3.3), which completes the proof of Theorem 3.11. ∎
3.3. Definition and examples of $G$-free independence
Now that we understand the combinatorics of moments for the $G$-free product, we define independence as follows:
Definition 3.16.
Let $(\mathcal{A},\phi)$ be a non-commutative probability space, let $G=(V,E)$ be a graph and let $(\mathcal{A}_{v})_{v\in V}$ be *-subalgebras. We say that $(\mathcal{A}_{v})_{v\in V}$ are $G$-freely independent if for every $k\in\mathbb{N}$, for every labelling $\ell:[k]\to V$, and for all $a_{j}\in B(\mathcal{H}_{\ell(j)})$ for $j=1$, …, $k$, we have
(3.6) $\phi(a_{1}\dots a_{k})=\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[a_{j}:j\in B].$
In other words, $G$-free independence means by definition that the algebras $\mathcal{A}_{v}$ have joint moments satisfying the conclusion of Theorem 3.11, or equivalently, they agree with the moments of operators on the $G$-free product Hilbert space (Definition 3.3) obtained from the GNS representations $L^{2}(\mathcal{A}_{v},\phi|_{\mathcal{A}_{v}})$ for $v\in V$.
Although for general $G$ we do not know how to describe $G$-free independence using simple condition on moments or cumulants, such as vanishing of certain mixed moments or some product formula, such conditions can be given for many different examples. How Theorem 3.11 relates to the moment conditions for boolean, monotone, and free independence has been discussed in depth in [10, §4.6, §7.3], so here let us focus on BM independence.
BM independence, defined by the third author [29], uses a poset to specify a mixture of boolean and monotone independence. In our terminology, the digraph $G=(V,E)$ is a poset if the adjacency relation $E\subseteq V\times V$ is strict partial order, specifically:
  • If $v\rightsquigarrow w$, then $w\not\rightsquigarrow v$.
  • If $v_{1}\rightsquigarrow v_{2}$ and $v_{2}\rightsquigarrow v_{3}$, then $v_{1}\rightsquigarrow v_{3}$.
[…]
Definition 3.17.
Given a poset $G=(V,E)=(V,\prec)$, we say that a family $(\mathcal{A}_{v})_{v\in V}$ in a non-commutative probability space $(\mathcal{A},\phi)$ is BM-independent if the following conditions hold for $v_{1},\dots,v_{n}\in V$ and $a_{j}\in\mathcal{A}_{\ell(j)}$ for $j=1$, …$n$:
BM1:
If $v_{j-1}\prec v_{j}\succ v_{j+1}$ or $v_{j-1}\prec v_{j}\nsim v_{j+1}$ or $v_{j-1}\nsim v_{j}\succ v_{j+1}$, then
(3.7) $\phi(a_{1}\dots a_{n})=\phi(a_{j})\phi(a_{1}\dots a_{j-1}a_{j+1}\dots a_{n})$
BM2:
If $v_{1}\succ\dots\succ v_{k}\nsim v_{k+1}\nsim\dots\nsim v_{\ell}\prec v_{\ell+1}\prec\dots\prec v_{n}$ for some $1\leq k\leq\ell\leq n$, then
(3.8) $\phi(a_{1}\ldots a_{n})=\prod_{j=1}^{n}\phi(a_{j}).$
The conditions above BM1 and BM2 allow one to compute all joint moments $\phi(a_{1}\cdots a_{n})$ of bm-independent random variables $a_{1},\ldots,a_{n}$ by [29, Lemmas 2.3,2.4] and an algorithm to evaluate joint moments using these conditions is given in [20, Remark 2.3].
We will show that the definition of BM independence in [28] agrees with our more general definition of $G$-independence when $G$ represents a poset. It will be useful first to observe the following alternative description of $\mathcal{NC}_{k}(\ell,G)$ when $G$ is a poset.
Definition 3.18 ([20, Definitions 3.8]).
Let $(V,\preceq)$ be a poset, let $\ell:[k]\to V$ be a labeling, and let $\pi\in\mathcal{NC}_{k}(\ell)$. We say that $\ell$ establishes strict BM order on $\pi$ if $B\prec B^{\prime}$ in $\pi$ implies that $\ell(B)\prec\ell(B^{\prime})$ in $V$.
Observation 3.19.
Let $G=(V,E)=(V,\prec)$ be a strict partial order, which we also view as a digraph. Let $\ell:[k]\to V$ and let $\pi\in\mathcal{NC}_{k}(\ell)$. Then the following are equivalent:
  1. (1)
    $\ell$ establishes strict BM order on $\pi$.
  2. (2)
    $\ell$ defines a digraph homomorphism $F(\pi)\to G$.
  3. (3)
    $\ell$ defines a strict poset homomorphism $\mathrm{F}(\pi)\to(V,\prec)$.
Here in (3), we view $\mathrm{F}(\pi)$ as a poset by taking the transitive closure of the edge relation. Moreover, a strict poset homomorphism by definition is a map that preserves strict inequality $\prec$.
Proof.
(1) $\iff$ (3) is immediate from the definitions. Moreover, (2) $\iff$ (3) is immediate from the definitions and transitivity of $\prec$. ∎
Now we prove the equivalence of two definitions of independence given by a finite poset.
Proposition 3.20.
Let $G=(V,E)$ be a digraph such that $E=\prec$ defines a strict partial order. Let $(\mathcal{A},\phi)$ be a non-commutative probability space and let $(\mathcal{A}_{v})_{v\in V}$ be *-subalgebras. Then $(\mathcal{A}_{v})_{v\in V}$ are BM independent in the sense of Definition 3.17 if and only if they are $G$-freely independent in the sense of Definition 3.16.
Proof.
First, suppose that Definition 3.16 holds. To check BM1, suppose $v_{j-1}\not\succeq v_{j}\not\preceq v_{j+1}$. Let us evaluate $\phi(a_{1}\dots a_{n})$ using (3.6) and show that it agrees with $\phi(a_{j})\phi(a_{1}\dots a_{j-1}a_{j+1}\dots a_{n})$.
We claim that for every partition $\pi$ appearing in (3.6), $\{j\}$ must be a singleton in $\pi$. Recall that a partition $\pi$ appears in (3.6) if and only if $\pi$ is consistently labelled by $\ell$ and the labelling defines a digraph homomorphism from $\mathrm{F}(\pi)$ to $G$, or equivalently it defines a strict poset homomorphism, that is, $B\prec B^{\prime}$ in $\pi$ implies that $\ell(B)\prec\ell(B^{\prime})$. Now let $B$ be the block containing $j$. Suppose for contradiction that there is some $i<j$ in $B$. Since $\ell(j-1)\neq\ell(j)$, we see that $i\neq j-1$ and the block containing $j-1$ is nested immediately inside $B$, and so we would need $\ell(j)\prec\ell(j-1)$, but this contradicts our assumption that $\ell(j-1)\prec\ell(j)$ or $\ell(j-1)\nsim\ell(j)$. Similarly, if we assume for contradiction that there is some $i>j$ in $B$, then we obtain a contradiction by a symmetrical argument since the block of $j+1$ would be nested immediately inside $B$.
Since $\pi$ has a singleton block at $j$, we obtain a non-crossing partition $\pi^{\prime}=\pi\setminus\{\{j\}\}$ of $[n]\setminus\{j\}$. Note that $\pi^{\prime}$ is compatible with $\ell^{\prime}=\ell_{[n]\setminus\{j\}}$ and $G$. Conversely, we claim that every partition $\pi^{\prime}$ compatible with $\ell^{\prime}$ and $G$ arises in the way, or equivalently, for every such $\pi^{\prime}$, the partition $\pi^{\prime}\cup\{\{j\}\}$ of $[n]$ is compatible with $\ell$ and $G$. To this end, we must consider some blocks $B_{1}$ and $B_{2}$ in $\pi$ with $B_{2}$ immediately nested inside $B_{1}$. Since $\pi^{\prime}$ is already compatible with $\ell^{\prime}$ and $G$, the only case to check is when $B_{2}=\{j\}$. Note that either $\ell(j-1)\prec\ell(j)$ or $\ell(j+1)\prec\ell(j)$. Suppose that $\ell(j-1)\prec\ell(j)$.
  • If $B_{1}$ contains $\ell(j-1)$, then $\ell(B_{1})=\ell(j-1)\prec\ell(j)=\ell(B_{2})$, so we are done.
  • If $B_{1}$ does not contain $\ell(j-1)$, then the block $B_{3}$ containing $\ell(j-1)$ is nested inside $B_{1}$, and hence $\ell(B_{1})\prec\ell(B_{3})=\ell(j-1)\prec\ell(j)=\ell(B_{2})$, so again we are done.
In the case where $\ell(j+1)\prec\ell(j)$, the argument is symmetrical.
Therefore, we obtain that
$\phi(a_{1}\dots a_{n})$ $=\sum_{\pi\in\mathcal{NC}_{n}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[a_{i}:i\in B]$
$=K_{\operatorname{bool},1}[a_{j}]\sum_{\pi^{\prime}\in\mathcal{NC}_{[n]\setminus\{j\}}(\ell^{\prime},G)}K_{\operatorname{bool},|B|}[a_{i}:i\in B],$
$=\phi(a_{j})\phi(a_{1}\dots a_{j-1}a_{j+1}\dots a_{n}),$
where $\mathcal{NC}_{[n]\setminus\{j\}}(\ell^{\prime},G)$ denotes the set of non-crossing partitions of $[n]\setminus\{j\}$ that are compatible with $G$ and $\ell^{\prime}$.
Next to check BM2, suppose that $v_{1}\succ\dots\succ v_{k}\nsim v_{k+1}\nsim\dots\nsim v_{\ell}\prec v_{\ell+1}\prec\dots\prec v_{n}$ for some $1\leq k\leq\ell\leq n$. Let $\pi$ be a partition compatible with $G$ and $\ell$. We claim that $\pi$ consist entirely of singletons. Suppose for contradiction that $i$ and $j$ are in the same block $B_{1}$ and $i<j$. Since $i<j$, we must have either $i<\ell$ or $j>k$. Suppose that $i\leq\ell$. Let $B_{2}$ be the block containing $i+1$. Then $B_{1}\neq B_{2}$ since our assumptions on $v_{1}$, …, $v_{n}$ implies that consecutive indices have distinct labels. Since $i<\ell$, we have that $\ell(i)\not\preceq\ell(i+1)$ by our assumptions on $v_{1}$, …, $v_{n}$, and this contradicts the condition $\ell(B_{1})\prec\ell(B_{2})$ needed for $\pi$ to be compatible with $\ell$ and $G$. If $j>k$, we obtain a contradiction by a symmetrical argument. Thus, the only possibility is that $\pi$ consistents of singletons, and therefore (3.6) reduces to $\pi(a_{1}\dots a_{n})=K_{\operatorname{bool},1}[a_{1}]\dots K_{\operatorname{bool},1}[a_{n}]=\phi(a_{1})\dots\phi(a_{n})$.
Therefore, we have shown that Definition 3.16 implies Definition 3.17. Conversely, suppose that Definition 3.17 holds. Let $\psi_{v}=\phi|_{\mathcal{A}_{v}}$. Construct another probability space $(\mathcal{B},\psi)$ as the $G$-free product of $(\mathcal{A}_{v},\psi_{v})$, and let $\mathcal{B}_{v}$ be the image of $\mathcal{A}_{v}$ in $\mathcal{B}$. Then the $\mathcal{B}_{v}$’s are $G$-freely independent by Theorem 3.11. Therefore, also the $\mathcal{B}_{v}$’s are BM-independent by the preceding argument. Recall by [29, Lemmas 2.3,2.4] that BM-independence uniquely determines the joint moments of elements from the different algebras. Since the $\mathcal{A}_{v}$’s and the $\mathcal{B}_{v}$’s are both BM-independent, the joint moments of elements $a_{j}\in\mathcal{A}_{v(j)}$ for $j=1$, …, $n$ viewed inside $\mathcal{A}$ must be the same as their joint moments when viewed inside $\mathcal{B}$. Thus, since the $\mathcal{B}_{v}$’s are $G$-freely independent in $\mathcal{B}$, it follows that the $\mathcal{A}_{v}$’s are $G$-freely independent in $\mathcal{A}$. ∎
3.4. Relationship with general tree independence
Now let us explain the relationship with tree independence from [10] and [9], which is necessary since we will use results from [9] later on. Note that in [10], the construction was done in the $\mathcal{B}$-valued setting
Let $\mathcal{T}_{\operatorname{free},n}$ be the rooted tree described as follows. The vertices are the alternating strings on the alphabet $[n]$, including the empty string. The empty string is the root vertex of $\mathcal{T}_{\operatorname{free},n}$, for each vertex $j_{m}\dots j_{1}$ in the tree, its children are the vertices $jj_{m}\dots j_{1}$ for $j\neq j_{m}$.
Let $\mathcal{T}$ be a connected subtree of $\mathcal{T}_{\operatorname{free},n}$, and let $(\mathcal{H}_{1},\xi_{1})$, …, $(\mathcal{H}_{n},\xi_{n})$ be pointed Hilbert spaces. Then define $\amalg_{\mathcal{T}}[(\mathcal{H}_{1},\xi_{1}),\dots,(\mathcal{H}_{n},\xi_{n})]$ as the pair $(\mathcal{H},\xi)$ where
(3.9) $\mathcal{H}=\bigoplus_{m\geq 0}\bigoplus_{j_{m}\dots j_{1}\in\mathcal{T}}\mathcal{H}_{j_{m}}^{\circ}\otimes\dots\otimes\mathcal{H}_{j_{1}}^{\circ}.$
This is a generalization of the construction we already explained for digraphs. Indeed, if $G=(V,E)$ is a digraph on vertex set $V=\{1,\dots,n\}$, then let
$\operatorname{Walk}(G)=\{j_{m}\dots j_{1}:m\geq 0,j_{m}\leftsquigarrow j_{m-1}\leftsquigarrow\dots\leftsquigarrow j_{1}\}=\{\varnothing\}\sqcup\bigsqcup_{m\geq 0}E_{m}^{\dagger}$
Then taking $\mathcal{T}=\operatorname{Walk}(G)$ in the $\mathcal{T}$-free product (3.9) will reduce to the $G$-free product of (3.1). See also [10, Definition 3.18].
In the general setting of $\mathcal{T}$-free products, the inclusion maps $B(\mathcal{H}_{j})\to B(\mathcal{H})$ are given as follows. Let
$S_{j}$ =$=\{j_{m}\dots j_{1}\in\mathcal{T}\text{ such that }jj_{m}\dots j_{1}\in\mathcal{T}\}$
$S_{j}^{\prime}$ =$=\{j_{m}\dots j_{1}\in\mathcal{T}\text{ such that }j\neq j_{m}\text{ and }jj_{m}\dots j_{1}\not\in\mathcal{T}\}.$
Note that in the case $\mathcal{T}=\operatorname{Walk}(G)$, then $S_{j}$ is the set of reverse paths such that the leftmost vertex $j_{m}\rightsquigarrow j$, and $S_{j}^{\prime}$ is the set of reverse paths such that $j_{m}\neq j$ and $j_{m}$ is not $\rightsquigarrow j$. Thus, the generalization of $\mathcal{H}_{\rightsquigarrow v}$ and $\mathcal{H}_{\perp v}$ are respectively
$\mathcal{H}_{S_{j}}$ $=\bigoplus_{j_{m}\dots j_{1}\in S_{j}}\mathcal{H}_{j_{m}}^{\circ}\otimes\dots\otimes\mathcal{H}_{j_{1}},$
$\mathcal{H}_{S_{j}^{\prime}}$ $=\bigoplus_{j_{m}\dots j_{1}\in S_{j}}\mathcal{H}_{j_{m}}^{\circ}\otimes\dots\otimes\mathcal{H}_{j_{1}}.$
Then just as in (3.2), we have a decomposition
$\mathcal{H}\cong[\mathcal{H}_{j}\otimes\mathcal{H}_{S_{j}}]\oplus\mathcal{H}_{S_{j}^{\prime}},$
and define the map $\iota_{j}:B(\mathcal{H}_{j})\to B(\mathcal{H})$ by
$\iota_{j}(a)=[a\otimes\operatorname{id}_{\mathcal{H}_{S_{j}}}]\oplus 0_{\mathcal{H}_{S_{j}^{\prime}}}.$
In the case $\mathcal{T}=\operatorname{Walk}(G)$, this reduces to Definition 3.4.
Next, we turn to the generalization of Theorem 3.11. This will again express $\langle\xi,\iota_{\ell(1)}(a_{1})\dots\iota_{\ell(k)}(a_{k})\xi\rangle$, where $a_{j}\in B(\mathcal{H}_{\ell(j)})$ and $\ell(1)$, …, $\ell(k)$ is alternating, though a sum of boolean cumulants indexed by partitions compatible with the given tree $\mathcal{T}$. Compatibility is described as follows.
Given a labelling $\ell:[k]\to[n]$ and a compatible partition $\pi$ in the sense of Definition 3.9, we say that $\pi$ and $\ell$ are compatible with $\mathcal{T}$ if the following condition holds: For each block $B\in\pi$, let $B_{0}\rightsquigarrow B_{1}\rightsquigarrow\dots\rightsquigarrow B_{m}=B$ be the unique path from a minimal (exterior) block up to $B$, in the nesting forest $\mathrm{F}(\pi)$. Then for every block $B$, we have $\ell(B_{m})\dots\ell(B_{0})\in\mathcal{T}$.
Another interpretation of this statement is as follows. As in [10, Definition 4.15], we can make $\mathrm{F}(\pi)$ into a tree $\operatorname{graph}(\pi)$ by adding a new vertex $\varnothing$, which will be the root and to which all the minimal (exterior) blocks in $\mathrm{F}(\pi)$ will be attached as children. This is analogous to the way that $\operatorname{Walk}(G)$ has the empty path $\varnothing$ added as the root vertex. Then compatibility of $\pi$, $\ell$, and $\mathcal{T}$ means precisely that there is a digraph homomorphism $\phi:\operatorname{graph}(\pi)\to\mathcal{T}$ preserving the root, such that for every block $B$, the first letter of $\phi(B)$ is $\ell(B)$; see also [10, Remark 4.20]. In the case where $\mathcal{T}=\operatorname{Walk}(G)$, then after deleting the root vertex, we get a digraph homomorphism from $\mathrm{F}(\pi)$ to $\operatorname{Walk}(G)\setminus\{\varnothing\}$. Now $\operatorname{Walk}(G)\setminus\{\varnothing\}$ is a union of $n$ branches, each branch representing the paths starting at a vertex $v\in[n]$; this construction is like the universal cover of a digraph, except that the paths are one-directional. Just like in the case of the universal cover, homomorphisms from a tree $\mathrm{F}(\pi)$ into $\operatorname{Walk}(G)$ correspond to homomorphisms $\mathrm{F}(\pi)\to G$. Thus, homomorphisms from $\operatorname{graph}(\pi)$ into $\mathcal{T}$ as in [10, Remark 4.20] reduce in the case of $\mathcal{T}=\operatorname{Walk}(G)$ to homomorphisms from $\mathrm{F}(\pi)$ into $G$ as in Definition 3.10.
4. Convolution and limit theorems
4.1. The compactly supported case
Given a digraph $G=(V,E)$ and compactly supported measures $(\mu_{v})_{v\in V}$, we define the $G$-free convolution $\boxplus_{G}((\mu_{v})_{v\in V}$ as follows. Let $(\mathcal{A}_{v},\phi_{v})$ be a non-commutative probability space and $x_{v}\in\mathcal{A}_{v}$ self-adjoint such that the spectral distribution of $x_{v}$ with respect to $\phi_{v}$ is $\mu_{v}$. Let $(\mathcal{A},\phi)$ be the $G$-free product of $(\mathcal{A}_{v},\phi_{v})_{v\in V}$ and let $\iota_{v}:\mathcal{A}_{v}\to\mathcal{A}$ the corresponding inclusion. Then $\boxplus_{G}((\mu_{v})_{v\in V})$ is defined to be the spectral distribution of $\sum_{v\in V}\iota_{v}(x_{v})$.
For this to be well-defined, one should verify that the specific choice of $(\mathcal{A}_{v},\phi_{v})$ and $x_{v}$ does not affect the final result, so long as $x_{v}$ has the distribution $\mu_{v}$. Since $\sum_{v\in V}\iota_{v}(x_{v})$ is a bounded operator, its spectral distribution is uniquely determined by its moments. Thus, it suffices to show that the moments of $\sum_{v\in V}\iota_{v}(x_{v})$ are uniquely determined by the moments of $x_{v}$. This will follow from the next result, where we compute the moments of $x$ using Theorem 3.11.
Lemma 4.1.
Let $G=(V,E)$ be a digraph. Let $(\mathcal{A},\phi)$ be the $G$-free product of non-commutative probability spaces $(\mathcal{A}_{v},\phi_{v})$. Let $x_{v}\in\mathcal{A}_{v}$ be self-adjoint. Let $x=\sum_{v\in V}\iota_{v}(x_{v})$. Then for $k\in\mathbb{N}$,
$\phi(x^{k})=\sum_{\ell:[k]\to V}\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(x_{\ell(B)}).$
Here $\ell$ is required to be constant on each block $B$, and so $\ell(B)$ denotes the constant value on that block. Moreover, $\kappa_{\operatorname{bool},|B|}(x_{\ell(B)})$ denotes the $|B|$th boolean cumulant, that is, $\kappa_{\operatorname{bool},|B|}(x_{\ell(B)})=K_{\operatorname{bool},|B|}(x_{\ell(B)},\dots,x_{\ell(B)})$.
Proof.
Using multilinearity,
$\phi(x^{k})=\phi\left(\left(\sum_{v\in V}\iota_{v}(x_{v})\right)^{k}\right)=\sum_{\ell:[k]\to V}\phi(\iota_{\ell(1)}(x_{\ell(1)})\dots\iota_{\ell(k)}(x_{\ell(k)})).$
By Theorem 3.11, this equals
$\sum_{\ell:[k]\to V}\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}K_{\operatorname{bool},|B|}[x_{\ell(j)}:j\in B].$
Now $\ell$ must be constant on each block $B$ in the above expression and hence we can write $K_{\operatorname{bool},|B|}[x_{\ell(j)}:j\in B]$ equivalently as the $|B|$th cumulant of $x_{\ell(B)}$. ∎
For our limit theorems, we focus on repeated convolutions of same measure. For simplicity of notation, we denote by $\boxplus_{G}(\mu)$ the convolution of $(\mu_{v})_{v\in V}$ where all the $\mu_{v}$’s are equal to $\mu$. The previous lemma implies the following.
Lemma 4.2.
Let $\mu\in\mathcal{P}(\mathbb{R})$ be compactly supported. Let $G=(V,E)$ be a finite digraph. Then we have
$m_{k}(\boxplus_{G}(\mu))=\sum_{\pi\in\mathcal{NC}_{k}}|\operatorname{Hom}(\mathrm{F}(\pi),G)|\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
Here $m_{k}$ denotes the $k$th moment of a measure and $\kappa_{\operatorname{bool},k}(\mu)$ denotes the $k$th boolean cumulant.
Proof.
Let $(\mathcal{A}_{v},\phi_{v})$ be non-commutative probability spaces and $x_{v}\in\mathcal{A}_{v}$ self-adjoint with distribution $\mu$. Let $x=\sum_{v\in V}\iota_{v}(x_{v})$ as in the previous lemma. By the previous lemma,
$m_{k}(\boxplus_{G}(\mu))=\phi(x^{k})$ $=\sum_{\ell:[k]\to V}\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(x_{\ell(B)})$
$=\sum_{\ell:[k]\to V}\sum_{\pi\in\mathcal{NC}_{k}(\ell,G))}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
Now we exchange the order of summation over $\ell$ and $\pi$ to get
$m_{k}(\boxplus_{G}(\mu))$ $=\sum_{\pi\in\mathcal{NC}_{k}}\sum_{\begin{subarray}{c}\ell:[k]\to V\\ \pi\in\mathcal{NC}_{k}(\ell,G)\end{subarray}}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu)$
$=\sum_{\pi\in\mathcal{NC}_{k}}|\{\ell:[k]\to V,\pi\in\mathcal{NC}_{k}(\ell,G)\}|\,\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
Recall that $\pi\in\mathcal{NC}_{k}(G,\ell)$ if and only if $\ell$ defines a digraph homomorphism $\mathrm{F}(\pi)\to G$. Therefore,
|$|\{\ell:[k]\to V\\ \pi\in\mathcal{NC}_{k}(\ell,G)\}|=|\operatorname{Hom}(\mathrm{F}(\pi),G)|,$
which establishes the desired formula. ∎
The next lemma is the first step of Theorem 1.1. In fact, it is a special case of the theorem when $\mu_{n}=\mu^{\uplus 1/|V_{n}|}$ and $\mu$ is compactly supported.
Lemma 4.3.
Suppose that $G_{n}=(V_{n},E_{n})$ is a sequence of digraphs such that for every finite tree $G^{\prime}=(V^{\prime},E^{\prime})$,
$\lim_{n\to\infty}\frac{1}{|V_{n}|^{|V^{\prime}|}}|\operatorname{Hom}(G^{\prime},G_{n})|=\beta_{G^{\prime}}.$
For a forest $G^{\prime}$ which is the disjoint union of rooted trees $G_{1}^{\prime}$, …, $G_{k}^{\prime}$, let us write
$\beta_{G^{\prime}}=\beta_{G_{1}^{\prime}}\dots\beta_{G_{k}^{\prime}}.$
Then for every compactly supported measure $\mu$, we have
(4.1) $\lim_{n\to\infty}m_{k}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))=\sum_{\pi\in\mathcal{NC}_{k}}\beta_{\mathrm{F}(\pi)}\kappa_{\operatorname{bool},\pi}(\mu).$
Moreover, denoting by $\operatorname{rad}(\mu)$ the radius of the support of the measure $\mu$, we have $\operatorname{rad}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))\leq 4\operatorname{rad}(\mu))$. Hence, $\lim_{n\to\infty}\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|})$ exists in $\mathcal{P}(\mathbb{R})$.
Proof.
First, note that if $G^{\prime}$ is a forest which is a disjoint union of rooted trees $G_{1}^{\prime}$, …, $G_{k}^{\prime}$, then a digraph homomorphism from $G^{\prime}\to G_{n}$ is equivalent to a $k$-tuple of digraph homomorphisms $G_{i}^{\prime}\to G_{n}$ for $i=1$, …, $k$, and thus
$|\operatorname{Hom}(G^{\prime},G_{n})|=|\operatorname{Hom}(G_{1}^{\prime},G_{n})|\dots|\operatorname{Hom}(G_{k}^{\prime},G_{n})|.$
Moreover, since $|V^{\prime}|=|V_{1}^{\prime}|+\dots+|V_{k}^{\prime}|$, we have
$\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\frac{|\operatorname{Hom}(G_{1}^{\prime},G_{n})|}{|V_{n}|^{|V_{1}^{\prime}|}}\dots\frac{|\operatorname{Hom}(G_{k}^{\prime},G_{n})|}{|V_{n}|^{|V_{k}^{\prime}|}}.$
Hence,
$\lim_{n\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\beta_{G_{1}^{\prime}}\dots\beta_{G_{k}^{\prime}}=\beta_{G}.$
In other words, the hypothesis that we assumed true when $G^{\prime}$ is a tree extends automatically to the case when $G^{\prime}$ is a forest.
By the previous lemma,
$m_{k}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))=\sum_{\pi\in\mathcal{NC}_{k}}|\operatorname{Hom}(\mathrm{F}(\pi),G)|\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
By definition of the boolean convolution powers, $\kappa_{\operatorname{bool},|B|}(\mu^{\uplus 1/|V_{n}|})=(1/|V_{n}|)\kappa_{\operatorname{bool},|B|}(\mu)$. Thus,
$m_{k}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))$ $=\sum_{\pi\in\mathcal{NC}_{k}}|\operatorname{Hom}(\mathrm{F}(\pi),G)|\prod_{B\in\pi}\frac{1}{|V_{n}|}\kappa_{\operatorname{bool},|B|}(\mu)$
$=\sum_{\pi\in\mathcal{NC}_{k}}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
Here $|\pi|$ is the number of blocks in $\pi$, which is the same as the number of vertices in $\mathrm{F}(\pi)$. By the foregoing argument, for each $\pi$, we have
$\lim_{n\to\infty}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}=\beta_{\mathrm{F}(\pi)}.$
Therefore, (4.1) holds.
Finally, we prove our estimate on the support radius of $\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|})$. First, by Lemma 3.8, one can see that
$|\kappa_{\operatorname{bool},k}(\mu)|=|K_{\operatorname{bool},k}(x_{v},\dots,x_{v})|\leq\lVert x_{v}\rVert^{k}=\operatorname{rad}(\mu)^{k},$
where $x_{v}$ is the operator of multiplication by $x$ in $\mathcal{A}_{v}=L^{\infty}(\mathbb{R},\mu)$. Thus, we estimate
$|m_{k}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))|$ $\leq\sum_{\pi\in\mathcal{NC}_{k}}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}\prod_{B\in\pi}|\kappa_{\operatorname{bool},|B|}(\mu^{\uplus 1/|V_{n}|})|$
$\leq\sum_{\pi\in\mathcal{NC}_{k}}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}\prod_{B\in\pi}\operatorname{rad}(\mu)^{|B|}$
$=\sum_{\pi\in\mathcal{NC}_{k}}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}\operatorname{rad}(\mu)^{k}.$
Since homomorphisms are functions from $V^{\prime}$ to $V_{n}$, we have
$\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G)|}{|V_{n}|^{|\pi|}}\leq 1.$
Moreover, the number of non-crossing partitions of $[k]$ is the $k$th Catalan number $C_{k}$, which satisfies $C_{k}\leq 4^{k}$. Hence, overall $|m_{k}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))|\leq 4^{k}\operatorname{rad}(\mu)^{k}$. Because this holds for all $k$, we conclude that $\operatorname{rad}(\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|}))\leq 4\operatorname{rad}(\mu)$.
Thus, the support radius of $\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|})$ is uniformly bounded for all $n$, and hence convergence in moments for this sequence is equivalent to convergence in $\mathcal{P}(\mathbb{R})$. This concludes the final claim of the lemma. ∎
4.2. The general case
In order to define the $G$-free convolution for probability measures $(\mu_{v})_{v\in V}$ which do not necessarily have bounded support, we first want to express the convolution operation using complex-analytic transforms. For a probability measure $\mu$ on $\mathbb{R}$, write
[…]
and
$K_{\mu}(z)=z-1/G_{\mu}(z).$
Unfortunately, one must be careful to distinguish this notation from the notation for cumulants. If $\mu$ is compactly supported, then the $K$-transform is related to boolean cumulants by the formula
$K_{\mu}(z)=\sum_{k=1}^{\infty}z^{k-1}\kappa_{\operatorname{bool},k}(\mu),$
where $\kappa_{\operatorname{bool},k}$ is the $k$th boolean cumulant of any random variable $X$ with distribution $\mu$. In the case of digraph independences, we have the following result.
Proposition 4.4 ([10, Proposition 6.9]).
Let $G$ be a digraph on the vertex set $[n]$. Let $\mu_{1}$, …, $\mu_{n}$ be compactly supported measures. For each $j$, let $\operatorname{Walk}(G,j)$ be the tree whose vertices are the empty path and all reverse paths that start at vertex $j$. Let
$\nu_{j}=\boxplus_{\operatorname{Walk}(G,j)}(\mu_{1},\dots,\mu_{n})$
be the convolution of $\mu_{1}$, …, $\mu_{n}$ with respect to the tree $\operatorname{Walk}(G,j)$ as in [10]. Then $\nu_{1},\dots,\nu_{n}$ satisfy the equations
(4.2) $K_{\nu_{i}}(z)=K_{\mu_{i}}\left(z-\sum_{j\leftsquigarrow i}K_{\nu_{j}}(z)\right),$
and we have
$K_{\boxplus_{G}(\mu_{1},\dots,\mu_{n})}(z)=\sum_{i=1}^{n}K_{\nu_{i}}(z).$
The system (4.2) is a fixed-point equation for $(K_{\nu_{1}}(z),\dots,K_{\nu_{n}}(z))$, which suggests a way to extend the definition of $\boxplus_{G}(\mu_{1},\dots,\mu_{n})$ to general probability measures $\mu_{1}$, …, $\mu_{n}$. It suffices to show that the solution to the fixed point equation exists, is unique, and depends continuously on the input measure. This was done in [9, Theorem 4.1] in the more general setting of tree convolutions, using the Earle-Hamilton theorem. This argument implies in particular that there is a unique $(\nu_{1},\dots,\nu_{n})$ satisfying (4.2). Therefore, the following definition is consistent:
Definition 4.5.
Let $G$ be a digraph on vertex set $[n]$. For $\mu_{1}$, …, $\mu_{n}\in\mathcal{P}(\mathbb{R})$, let $(\nu_{1},\dots,\nu_{n})$ satisfy (4.2). Then $\boxplus_{G}(\mu_{1},\dots,\mu_{n})$ is defined to be $\nu_{1}\uplus\dots\uplus\nu_{n}$. Moreover, in the case when the measures $\mu_{j}$ are the same, we write
$\boxplus_{G}(\mu)=\boxplus_{G}(\mu,\dots,\mu).$
Continuous dependence of the measures $\nu_{1}$, …, $\nu_{n}$ and consequently $\boxplus_{G}(\mu_{1},\dots,\mu_{n})$ upon the inputs $\mu_{1}$, …, $\mu_{n}$ also follows from [9, Theorem 4.1]. In fact, there is a stronger equicontinuity result [9, Theorem 6.2] that we will need for the proof of Theorem 1.1 in the case of measures with unbounded support. Here we will use the Lévy distance on $\mathcal{P}(\mathbb{R})$, given by
$d_{L}(\mu,\nu):=\inf\Bigl{\{}\epsilon>0:\mu((-\infty,x-\epsilon))-\epsilon\leq\nu((-\infty,x))\leq\mu(-\infty,x+\epsilon))+\epsilon)\text{ for all }x\in\mathbb{R}\Bigr{\}}.$
The distance $d_{L}$ makes $\mathcal{P}(\mathbb{R})$ into a complete metric space, and the induced topology is the same as the weak-* topology from viewing $\mathcal{P}(\mathbb{R})$ inside the dual of $C_{0}(\mathbb{R})$; see for instance [5, Theorem 6.8]. Here we state [9, Theorem 6.2] specialized to digraph convolutions.
Proposition 4.6 ([9, Theorem 6.2]).
Let $d_{L}$ be the Lévy distance on $\mathcal{P}(\mathbb{R})$. For every $Y\subseteq\mathcal{P}(\mathbb{R})$ compact and $\epsilon>0$, there exists $\delta>0$ such that for every digraph $G=(V,E)$ and every $\mu\in Y$ and $\nu\in\mathcal{P}(\mathbb{R})$,
$d_{L}(\mu,\nu)<\delta\implies d_{L}(\boxplus_{G}(\mu^{\uplus 1/|V|}),\boxplus_{G}(\nu^{\uplus 1/|V|}))<\epsilon.$
Now we can conclude the proof of the main theorem for probability measures with unbounded support.
Proof of Theorem 1.1.
Let $G_{n}=(V_{n},E_{n})$ be a sequence of digraphs such that $\lim_{n\to\infty}|V_{n}|=\infty$ and for every finite rooted tree $G^{\prime}=(V^{\prime},E^{\prime})$, the limit $\beta_{G^{\prime}}=\lim_{n\to\infty}|\operatorname{Hom}(G^{\prime},G)|/|V_{k}|^{|V^{\prime}|}$ exists. Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence of probability measures such that $\lim_{n\to\infty}\mu_{n}^{\uplus|V_{n}|}=\mu$, and write $\nu_{n}=\mu_{n}^{\uplus|V_{n}|}$. In order to show that $(\boxplus_{G_{k}}(\mu_{n}))_{n\in\mathbb{N}}$ is a Cauchy sequence in $d_{L}$, fix $\epsilon>0$. Then $Y=\{\nu_{n}:n\in\mathbb{N}\}\cup\{\mu\}$ is compact. By Proposition 4.6, there exists $\delta>0$ such that for all $\nu$, we have
$d_{L}(\nu,\nu_{n})<\delta\implies d_{L}(\boxplus_{G_{n}}(\nu^{\uplus 1/|V_{n}|}),\boxplus_{G_{n}}(\nu_{n}^{\uplus 1/|V_{n}|}))<\frac{\epsilon}{4},$
and the same holds with $\nu_{n}$ replaced by $\mu$. Let $\sigma_{R}=\mu([-R,R])^{-1}\mu|_{[-R,R]}$. By choosing $R$ sufficiently large, we can arrange that $d_{L}(\mu,\sigma_{R})<\delta$. For sufficiently large $n$, we also have $d_{L}(\nu_{n},\sigma_{R})<\delta$ as well. Hence,
$d_{L}(\boxplus_{G_{n}}(\mu_{n}),\boxplus_{G_{n}}(\sigma_{R}^{\uplus 1/|V_{n}|}))=d_{L}(\boxplus_{G_{n}})(\nu_{n}^{\uplus 1/|V_{n}|}),\boxplus_{G_{n}}(\sigma_{R}^{\uplus 1/|V_{n}|}))<\frac{\epsilon}{4}.$
By Lemma 4.3, $\sigma_{R}^{\prime}:=\lim_{n\to\infty}\boxplus_{G_{n}}(\sigma_{R}^{\uplus 1/|V_{n}|})$ exists, and therefore for sufficiently large $n$,
$d_{L}(\boxplus_{G_{n}}(\sigma_{R}^{\uplus 1/|V_{n}|}),\sigma_{R}^{\prime})<\frac{\epsilon}{4}.$
Using the triangle inequality, for sufficiently large $n$ and $m$,
$d_{L}(\boxplus_{G_{n}}(\mu_{n}),\boxplus_{G_{m}}(\mu_{m}))<\epsilon.$
Hence, $(\boxplus_{G_{n}}(\mu_{n}))_{n\in\mathbb{N}}$ is Cauchy in $d_{L}$ and hence converges to some limit $\mu^{\prime}$.
Similar reasoning shows that if $\epsilon$ and $\delta$ are as above and $d(\sigma_{R},\mu)<\delta$, then
$d_{L}(\mu^{\prime},\sigma_{R}^{\prime})<\epsilon.$
Hence,
$\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})=\mu^{\prime}=\lim_{R\to\infty}\sigma_{R}^{\prime}.$
Since $\sigma_{R}$ is the truncation of $\mu$, it only depends on $\mu$. Moreover, $\sigma_{R}^{\prime}$ given by Lemma 4.3 only depends on $\mu$ and the coefficients $\beta_{G^{\prime}}$. Therefore, $\mu^{\prime}$ only depends on $\mu$ and the coefficients $\beta_{G^{\prime}}$. ∎
5. Examples and applications
In this section, we describe several classes of examples to which Theorem 1.1 applies.
5.1. Continuum limit method
In [11, 12, 29, 20, 19, 21], certain limit theorems for BF and BM independence associated to cones were obtained using Riemann sum approximations. BM independence was defined using finite posets which were discretizations of a bounded region in a cone (these posets in our setting can be viewed as digraphs). In this section, we give a continuum limit method in a more general measure-theoretic constext and hence prove Proposition 1.2.
Let $(\Omega,\rho)$ be a complete probability measure space (for a concrete example, one could take $\Omega=[0,1]$ with Lebesgue measure). Let $\mathcal{E}\subseteq\Omega\times\Omega$ be measurable. We view $\Omega$ as a vertex set and $\mathcal{E}$ as an edge set, so $(\Omega,\mathcal{E})$ is a “measurable diagraph.”
Now fix a digraph $G^{\prime}=(V^{\prime},E^{\prime})$. Note that homomorphisms $(V^{\prime},E^{\prime})$ to $(\Omega,\mathcal{E})$ can be described as functions $\phi:V^{\prime}\to\Omega$ such that if $(v,w)\in E$, then $(\phi(v),\phi(w))\in\mathcal{E}$. Functions $V^{\prime}\to\Omega$ may be identified with the Cartesian product $\Omega^{\times V^{\prime}}$, and so we obtain
$\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}))=\{\omega\in\Omega^{\times V^{\prime}}:(v,w)\in E\implies(\omega_{v},\omega_{w})\in\mathcal{E}\}.$
Since $(\Omega^{\times V^{\prime}},\rho^{\times V^{\prime}})$ is a probability space, it makes sense to evaluate the measure of the space of homomorphisms:
$\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E})))=\int_{\Omega^{\times V^{\prime}}}\prod_{(v,w)\in E^{\prime}}\mathbf{1}_{\mathcal{E}}(\omega_{v},\omega_{w})\,d\rho^{\times V^{\prime}}(\omega),$
where $\omega=(\omega_{v})_{v\in V^{\prime}}\in\Omega^{\times V^{\prime}}$.
We will show that if $G_{n}$ is a sequence of graphs giving a discretization of $(\Omega,\mathcal{E})$, then the normalized count of homomorphisms from $G^{\prime}$ to $G_{n}$ as in Theorem 1.1 converges to $\beta_{G^{\prime}}:=\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E})))$; see Proposition 1.2 below.
We can relate a finite digraphs and measureable digraphs as follows. Let $G=(V,E)$ be a finite digraph. Let $(A_{v})_{v\in V}$ be a partition of $\Omega$ into measureable sets with $\rho(A_{v})=1/|V|$ for all $v\in V$. (For example, if $\Omega=[0,1]$ and $V=\{1,\dots,k\}$, we could take $A_{j}=[(j-1)/k,j/k)$.) Let
(5.1) ~$\tilde{\mathcal{E}}=\bigcup_{(v,w)\in E}A_{v}\times A_{w}\subseteq\Omega\times\Omega.$
Then we claim that
(5.2) $\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\tilde{\mathcal{E}})))=\frac{|\operatorname{Hom}(G^{\prime},G)|}{|V|^{|V^{\prime}|}}.$
To see this, suppose $\omega\in\Omega^{V^{\prime}}$ and note there is a unique $\phi:V^{\prime}\to V$ such that $\omega_{v}\in A_{\phi(v)}$ for each $v\in V$. Moreover, $\omega\in\operatorname{Hom}(G^{\prime},(\Omega,\tilde{\mathcal{E}}))$ if and only if $\phi\in\operatorname{Hom}(G^{\prime},G)$. Thus,
$\operatorname{Hom}(G^{\prime},(\Omega,\tilde{\mathcal{E}}))=\sum_{\phi\in\operatorname{Hom}(G^{\prime},G)}\prod_{v\in V^{\prime}}A_{\phi(v)},$
and since $\prod_{v\in V^{\prime}}A_{\phi(v)}$ has measure $1/|V|^{|V^{\prime}|}$, we obtain (5.2).
Observation 5.1.
Let $(\Omega,\rho)$ be a probability measure space. Let $\mathcal{E},\mathcal{E}^{\prime}\subseteq\Omega\times\Omega$ be measurable sets and $\mathcal{E}\Delta T$ their symmetric difference. Let $G^{\prime}=(V^{\prime},E^{\prime})$ be a finite digraph. Then
$|\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E})))$ $-\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}^{\prime})))|$
$\leq\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}))\Delta\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}^{\prime})))$
$\leq|E^{\prime}|\rho^{\times 2}(S\Delta T).$
Proof.
The first inequality is immediate. For the second, note that
$\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}))\Delta\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}^{\prime})))=\left\lVert\prod_{(v,w)\in E^{\prime}}\mathbf{1}_{\mathcal{E}}(v,w)-\prod_{(v,w)\in E^{\prime}}\mathbf{1}_{\mathcal{E}^{\prime}}(v,w)\right\rVert_{L^{1}(\rho^{\times V^{\prime}})}.$
We swap out each $\mathbf{1}_{\mathcal{E}}$ for $\mathbf{1}_{\mathcal{E}^{\prime}}$ one instance at a time. Each swap produces an error of at most $\lVert\mathbf{1}_{\mathcal{E}}-\mathbf{1}_{\mathcal{E}^{\prime}}\rVert_{L^{1}(\rho^{\times 2})}=\rho^{\times 2}(\mathcal{E}\Delta\mathcal{E}^{\prime})$ because the product of the other terms is zero or one. Overall there are $|E^{\prime}|$ swaps, and so the error is at most $|E^{\prime}|\rho^{\times 2}(\mathcal{E}\Delta\mathcal{E}^{\prime})$. ∎
Proof of Proposition 1.2.
The first claim follows from (5.2) and Observation 5.1, and the second claim follows from Theorem 1.1. ∎
5.2. Limit theorems for BM-independence associated to cones
One of the motivating examples for the continuum limit approach is the case of BM independences for cones studied by the second and third author in [28, 29, 20, 19, 21], and the similar results for BF independence of the third author with Kula [11, 12]. In this subsection, we will give a generalization of the BM limit theorems from [20] using Theorem 1.1 and Proposition 1.2 from this paper, as well as explain how the technique of [20] closely relate to the proofs given in this paper.
First, we recall some terminology relating to convex cones. We say $\Pi\subseteq\mathbb{R}^{d}$ is a convex cone if it is closed under addition and positive scalar multiples. We assume that $\Pi$ is closed and that it is salient, meaning that $\Pi\cap-\Pi=\{0\}$. In this case, the relation $\preceq$ on $\mathbb{R}^{d}$ defined by setting $\xi\preceq\eta$ if and only if $\eta-\xi\in\Pi$ is a partial order. We define the interval
$[\xi,\eta]=\{\rho\in\mathbb{R}^{d}:\xi\preceq\rho\preceq\eta\}$
(which is nonempty if and only if $\xi\leq\eta$). Examples of convex cones studied in [19] include the positive orthant $\mathbb{R}_{+}^{d}\subseteq\mathbb{R}^{d}$, the Lorentz light-cone
$\Lambda_{d}^{1}=\{(t;\mathbf{x})\in\mathbb{R}^{d+1}:t\geq\lVert x\rVert\},$
and the positive semidefinite matrices $M_{d}(\mathbb{R})_{+}$ which is a subset of the space of symmetric matrices $\operatorname{Sym}_{d}(\mathbb{R})\cong\mathbb{R}^{d(d+1)/2}$.
Given a salient closed convex cone $\Pi\subseteq\mathbb{R}^{d}$, one can obtain finite posets by considering $I_{\xi}=[0,\xi]\cap\mathbb{Z}^{d}$ for $\xi\in\Pi$. Then, as in §3.3, one can consider BM-independent random variables indexed by $I_{\xi}$, which by Proposition 3.20 is equivalent to $I_{\xi}$-independent variables where we view $I_{\xi}$ is a digraph. We will study the behavior of $\boxplus_{I_{\xi}}(\mu)$ as $\xi\xrightarrow{\Pi}\infty$, or as $\xi$ tends to infinity in the cone $\Pi$. Here we recall that if $f$ is a function on the cone $\Pi$, we say that $f(\xi)\to L$ as $\xi\xrightarrow{\Pi}\infty$, if for every $\epsilon>0$, there exists $\xi_{0}\in\Pi$ such that for all $\xi\succeq\xi_{0}$ we have $|f(\xi)-L|<\epsilon$. The meaning of $\xi\xrightarrow{\Pi}\infty$ in the specific cases of $\mathbb{R}_{+}^{d}$, $\Lambda_{d}^{1}$, and $M_{d}(\mathbb{R})_{+}$ is explained in [20, Definition 1.5].
By Theorem 1.1, we need to study the limit as $\xi\xrightarrow{\Pi}\infty$ of $|\operatorname{Hom}(G^{\prime},I_{\xi})|/|I_{\xi}|^{|V^{\prime}|}$ for a finite forest $G^{\prime}=(V^{\prime}E^{\prime})$. We remark that $\operatorname{Hom}(G^{\prime},I_{\xi})$ is equivalently the set of strict poset homomorphisms from $G^{\prime}$ to $I_{\xi}$ by Observation 3.19, and in the case that $G^{\prime}=\mathrm{F}(\pi)$ for some non-crossing partition $\pi$, this is exactly the set of $I_{\xi}$ labelings that establish strict BM order on $\pi$; this is denoted by $\operatorname{BMO}(\pi;\xi)$ in [20, Definition 4.1]. The limit of $|\operatorname{BMO}(\pi;\xi)|/|I_{\xi}|^{|\pi|}$ is described in [20, Theorem 4.4, Corollary 4.5]. We will explain the computation here in two steps, first applying the continuum limit method, and then computing the volume of the limiting set explicitly using the volume characteristic of [11].
Lemma 5.2.
Let $\Pi$ be one of the cones as above. Let $G^{\prime}=(V^{\prime},E^{\prime})$ be a finite forest with a rooted assigned in each component, viewed as a poset. Let
$\operatorname{Hom}(G^{\prime},[0,\xi])=\{\eta\in[0,\xi]^{\times V^{\prime}}:v\rightsquigarrow w\text{ in }G^{\prime}\implies\eta_{v}\prec\eta_{w}\text{ in }\Pi\}\subseteq(\mathbb{R}^{d})^{\times V^{\prime}}.$
Then
$\lim_{\xi\xrightarrow{\Pi}\infty}\left|\frac{|\operatorname{Hom}(G^{\prime},I_{\xi})|}{|I_{\xi}|^{|V^{\prime}|}}-\frac{\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))}{\operatorname{vol}([0,\xi])^{|V^{\prime}|}}\right|=0.$
The idea of this lemma is the same as Proposition 1.2, but here we do not have a fixed continuum limit, since the continuum object $\operatorname{Hom}(G^{\prime},[0,\xi])$ also depends on $\xi$. And of course, we are taking the limit as $\xi\xrightarrow{\Pi}\infty$ rather than only a limit as $N\to\infty$. Thus, we must proceed carefully to define the discretized set and estimate the symmetric difference. Here we will leave some details to the reader since [20] gives an alternative argument for the limit in 5.2.
Lemma 5.3.
Let $\Pi\subseteq\mathbb{R}^{d}$ be one of the cones above. For $\xi\in\mathbb{R}^{d}$, let $Q_{\eta}$ be the unit cube $\prod_{j=1}^{d}[\xi_{j}-1/2,\xi_{k}+1/2]$. Let
$A_{\xi}=\bigcup_{\eta\in I_{\xi}}Q_{\eta}$
Then
$\lim_{\xi\xrightarrow{\Pi}\infty}\frac{\operatorname{vol}(A_{\xi}\Delta[0,\xi])}{\operatorname{vol}([0,\xi])}=0.$
Similarly, let
$E_{\xi}$ =$=\{(\eta_{1},\eta_{2})\in\mathbb{Z}^{d}\times\mathbb{Z}^{d}:0\preceq\eta_{1}\prec\eta_{2}\preceq\xi\}$
$B_{\xi}$ $=\bigcup_{(\eta_{1},\eta_{2})\in E_{\xi}\cap\mathbb{Z}^{2d}}Q_{(\eta_{1},\eta_{2})}.$
Then
$\lim_{\xi\xrightarrow{\Pi}\infty}\frac{\operatorname{vol}(B_{\xi}\Delta E_{\xi})}{\operatorname{vol}([0,\xi])^{2}}=0.$
Proof.
For the first claim, we note that $A_{\xi}\,\Delta\,[0,\xi]$ is contained in the union of the cubes that intersect the boundary $\partial[0,\xi]$. Hence, in particular, letting $N_{\delta}(\partial[0,\xi])$ of a set be the closed $\delta$-neighborhood in the $\ell^{\infty}$ metric on $\mathbb{R}^{d}$, we have
$\operatorname{vol}(A_{\xi}\Delta[0,\xi])\leq\operatorname{vol}(N_{1/2}(\partial[0,\xi])),$
and so the claim reduces to proving that
$\lim_{\xi\xrightarrow{\Pi}\infty}\frac{\operatorname{vol}(N_{1/2}(\partial[0,\xi]))}{\operatorname{vol}[0,\xi]}.$
This can be proved by explicit estimates in each of the three cases of $\Pi$ under consideration here. We leave the details to the reader. For the second claim, one can similarly reduce to the showing that
$\lim_{\xi\xrightarrow{\Pi}\infty}\frac{\operatorname{vol}(N_{1/2}(\partial E_{\xi}))}{\operatorname{vol}[0,\xi]^{2}}=0,$
and then perform direct estimates for each case of $\Pi$. ∎
Lemma 5.2 follows from Lemma 5.3 by similar reasoning as we used in Observation 5.1.
It remains to compute the volume of $|\operatorname{Hom}(G^{\prime},[0,\xi])|$ appearing in Lemma 5.2. This computation drastically simplifies due to the special geometric structure of the cones under consideration, as shown by Kula and the third author in [11].
Proposition 5.4 (Volume characteristic[11]).
For each of the positive symmetric cones $\Pi$ we consider there exists a sequence $(\gamma_{n}(\Pi))_{n\geq 1}$ such that for any $\xi\in\Pi$ and any $n\in\mathbb{N}$
$\int_{\rho\in[0,\xi]}\operatorname{vol}([0,\rho])^{n-1}d(\rho)=\gamma_{n}(\Pi)\operatorname{vol}([0,\xi])^{n}.$
This allows for a recursive computation of the volume of $\operatorname{Hom}(G^{\prime},[0,\xi])$, as described in [20, Theorem 4.4, Corollary 4.5]. Here we express the result of the computation explicitly as a product rather than giving a recursive description as in [20].
Lemma 5.5.
For the cones $\Pi$ under consideration and $\xi\in\Pi$ and for a finite rooted forest $G^{\prime}=(V^{\prime},E^{\prime})$, we have
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))=\operatorname{vol}([0,\xi])^{|V^{\prime}|}\prod_{v\in V^{\prime}}\gamma_{k(v)},$
where $k(v)=|\{v\in V^{\prime}:v\succeq w\}|$.
Proof.
We proceed by induction on $V^{\prime}$. If $|V^{\prime}|=1$, then both sides are equal to $\operatorname{vol}[0,\xi]$.
Next, suppose that $|V^{\prime}|>1$, and suppose that $G^{\prime}$ has more than one connected component (here components are defined by forgetting the orientation of the edges). Write $G^{\prime}$ as the disjoint union of components $G_{1}^{\prime}$, …, $G_{k}^{\prime}$. Then
$\operatorname{Hom}(G^{\prime},[0,\xi])\cong\operatorname{Hom}(G_{1}^{\prime},[0,\xi])\times\dots\times\operatorname{Hom}(G_{k}^{\prime},[0,\xi]).$
By applying the induction hypothesis to $G_{j}^{\prime}$, we get
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))=\prod_{j=1}^{k}\left(\operatorname{vol}([0,\xi])^{|V_{j}^{\prime}|}\prod_{v\in V_{j}^{\prime}}\gamma_{k(v)}\right)=\operatorname{vol}([0,\xi])^{|V^{\prime}|}\prod_{v\in V_{j}^{\prime}}\gamma_{k(v)}.$
Finally, suppose that $|V^{\prime}|>1$ and that $G^{\prime}$ has only one component, i.e. $G^{\prime}$ is a rooted tree. Let $r$ be the root vertex, let $v_{1}$, …, $v_{k}$ be the children of $r$, and let $G_{j}^{\prime}$ be the rooted subtree under $v_{j}^{\prime}$. Observe that
$\operatorname{Hom}(G^{\prime},[0,\xi])=\{(\eta,\eta_{1},\dots,\eta_{k}):\eta\in[0,\xi],\eta_{j}\in\operatorname{Hom}(G_{j}^{\prime},(\eta,\xi])\},$
where $(\eta,\xi]=\{\zeta:\eta\prec\zeta\preceq\xi\}$; this follows by first choosing the point $\eta$ where the root $r$ is mapped and then restricting the homomorphism to each of the branches $G_{j}$. It follows from Fubini-Tonelli that
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))=\int_{[0,\xi]}\prod_{j=1}^{k}\operatorname{vol}(\operatorname{Hom}(G_{j}^{\prime},(\eta,\xi]))\,d\eta.$
By ignoring the boundary, we can use $[\eta,\xi]$ instead of $(\eta,\xi]$. Now perform the change of variables $\eta\mapsto\xi-\eta$ to obtain
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))=\int_{[0,\xi]}\prod_{j=1}^{k}\operatorname{vol}(\operatorname{Hom}(G_{j}^{\prime},[0,\eta]))\,d\eta.$
Applying the induction hypothesis to $G_{j}^{\prime}$, we obtain
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))$ $=\int_{[0,\xi]}\prod_{j=1}^{k}\left(\operatorname{vol}([0,\eta])^{|V_{j}^{\prime}|}\prod_{v\in V_{j}^{\prime}}\gamma_{k(v)}\right)\,d\eta$
$=\int_{[0,\xi]}\operatorname{vol}([0,\eta])^{|V^{\prime}|-1}\,d\eta\cdot\prod_{v\in V^{\prime}\setminus\{r\}}\gamma_{k(v)}.$
By Proposition 5.4,
$\operatorname{vol}(\operatorname{Hom}(G^{\prime},[0,\xi]))$ $=\operatorname{vol}([0,\xi])^{|V^{\prime}|}\gamma_{|V^{\prime}|}\prod_{v\in V^{\prime}\setminus\{r\}}\gamma_{k(v)}$
$=\operatorname{vol}([0,\xi])^{|V^{\prime}|}\prod_{v\in V^{\prime}}\gamma_{k(v)}$
since $k(r)=|V^{\prime}|$. ∎
Remark 5.6.
Note that analogous computations in [20] are written in terms of the partition $\pi$ rather than forest $G^{\prime}$, and thus correspond to taking $G^{\prime}=\mathrm{F}(\pi)$. The case of several connected components $G_{1}^{\prime}$, …, $G_{k}^{\prime}$ corresponds to when $\pi$ is the disjoint union / concatenation of partitions $\pi_{1}$, …, $\pi_{k}$. Similarly, if $G^{\prime}$ has only one component and we look at the branches $G_{j}^{\prime}$, this corresponds to taking a partition $\pi$ with only one outer block $B$ and looking at the subpartitions $\pi_{1}$, …, $\pi_{k}$ in between consecutive elements of the block $B$.
Putting together Lemmas 5.2 and 5.5 with Theorem 1.1, we obtain the following result.
Theorem 5.7 (BM limit theorems for cones).
Let $\Pi$ be one of the cones $\mathbb{R}_{+}^{d}$, $\Lambda_{d}^{1}$, or $M_{d}(\mathbb{R})_{+}$, and let $I_{\xi}$ be as above. For a finite rooted forest $G^{\prime}=(V^{\prime},E^{\prime})$, we have
$\lim_{\xi\xrightarrow{\Pi}\infty}\frac{|\operatorname{Hom}(G^{\prime},I_{\xi})|}{|I_{\xi}|^{|V^{\prime}|}}=\prod_{v\in V^{\prime}}\gamma_{k(v)}.$
In particular, by Theorem 1.1, if $\mu_{\xi}$ is a family of probability measures such that $\lim_{\xi\xrightarrow{\Pi}\infty}\mu_{\xi}^{\uplus|I_{\xi}|}=\mu$, then $\widehat{\mu}=\lim_{\xi\xrightarrow{\Pi}\infty}\boxplus_{I_{\xi}}(\mu)$ exists. Moreover, in light of Lemma 4.3, if $\mu$ is compactly supported, then so is $\widehat{\mu}$, and
$m_{k}(\widehat{\mu})=\sum_{\pi\in\mathcal{NC}_{k}}\prod_{v\in\mathrm{F}(\pi)}\gamma_{k(v)}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu).$
Example 5.8 (Poisson limit theorem for BM independence).
The law of small numbers or the Poisson limit theorem for BM independence studied in [20] is a special case of Theorem 5.7. For the Poisson limit theorem, we must plug in for $\mu$ the boolean analog of the Poisson distribution, which turns out to be $\frac{1}{1+\lambda}\delta_{0}+\frac{\lambda}{1+\lambda}\delta_{1+\lambda}$. Indeed, one can show by direct computation of $K$-transforms that
$[(1-\lambda/n)\delta_{0}+(\lambda/n)\delta_{1}]^{\uplus n}\to\frac{1}{1+\lambda}\delta_{0}+\frac{\lambda}{1+\lambda}\delta_{1+\lambda};$
and that the boolean cumulants of $\frac{1}{1+\lambda}\delta_{0}+\frac{\lambda}{1+\lambda}\delta_{1+\lambda}$ are all equal to $\lambda$. By Theorem 5.7, if $\mu_{\xi}$ is some measure with $\mu_{\xi}^{\uplus|I_{\xi}|}\to\mu$, then we have $\boxplus_{I_{\xi}}(\mu_{\xi})\to\widehat{\mu}$ where
$m_{k}(\widehat{\mu})=\sum_{\pi\in\mathcal{NC}_{k}}\prod_{v\in\mathrm{F}(\pi)}\gamma_{k(v)}\prod_{B\in\pi}\lambda^{|\pi|}.$
This is the same result as [20, Theorem 4.4] up to some technical differences.
Specifically, [20, Theorem 4.4] allowed the convolution of several different measures, rather than only copies of the same measure. Our result Theorem 1.1 could similarly be generalized to consider $\boxplus_{G_{n}}(\mu_{n,1},\dots,\mu_{n,|V_{n}|})$ where $\lim_{n\to\infty}\sup_{j=1,\dots,n}d_{L}(\mu_{n,j},\mu)=0$ after generalizing [9, Theorem 6.2] to allow several different input measures. However, that is beyond the scope of this work.
Note also that the hypotheses and conclusion of [20, Theorem 4.4] use convergence of moments rather than weak-* convergence of measures, and neither type of convergence implies the other in general.
5.3. Iterated composition of digraphs
Another motivating case of the continuum limit method is the setting of iterated composition of digraphs studied in [10] (which of course also worked in the more general setting of tree independences).
First, we recall from [10, §5.5] the composition operation on digraphs. Let $\operatorname{Digraph}(n)$ be the set of directed graphs on $[n]$. Let $G\in\operatorname{Digraph}(k)$ and let $G_{j}\in\operatorname{Digraph}(n_{j})$ for $j=1$, …, $k$. Let $N=n_{1}+\dots+n_{k}$ and let $\iota_{j}:[n_{j}]\to[N]$ be the inclusion $\iota_{j}(i)=n_{1}+\dots+n_{j-1}+i$. Then $G(G_{1},\dots,G_{k})$ is the digraph $G^{\prime}$ on vertex set $[N]$ described by
$E^{\prime}=\{(\iota_{j}(v),\iota_{j}(w)):v\rightsquigarrow w\text{ in }G_{j}\}\cup\{(\iota_{i}(v),\iota_{j}(w)):i\rightsquigarrow j\text{ in }G,v\in V_{i},w\in V_{j}\}.$
In other words, we create disjoint copies of $G_{1}$, …, $G_{k}$, and then whenever $i\rightsquigarrow j$ in $G$ we add edges from every vertex in $G_{i}$ to every vertex in $G_{j}$. This composition operation defines a (symmetric) operad structure (see [13] for background on operads).
We focus here on iterated compositions of a fixed graph $G\in\operatorname{Digraph}(n)$. Define inductively $G^{\circ k}$ by $G^{\circ 1}=G$ and $G^{\circ(k+1)}=G(G^{\circ k},\dots,G^{\circ k})$. Limit theorems for such iterated compositions of the same digraph are given in [10] and [9]. The idea is essentially a continuum limit construction, where the limiting measure space is an infinite product, and the finite approximants are given by cylinder sets.
As motivation, let us describe the edge structure in $G^{\circ k}$, starting with $G^{\circ 2}$. The vertex set of $G^{\circ 2}$ is $[n]^{2}$, which we view as $[n]\times[n]$, where the first coordinate describes the position in the outer graph in the composition (i.e. which of the $n$ copies of $G$ you are in), and the second coordinate describes the position in the inner graph in the composition. Then $(i_{1},i_{2})\rightsquigarrow(j_{1},j_{2})$ if and only if either $i_{1}\rightsquigarrow i_{2}$ in the outer graph, or $i_{1}=i_{2}$ and $j_{1}\rightsquigarrow j_{2}$ in the inner graph. Similarly, the vertex set of $G^{\circ k}$ can be described as $[n]^{k}$ where the first coordinate corresponds to the outermost graph and the last coordinate corresponds to the innermost graph in the composition. To determine when $(i_{1},\dots,i_{k})\rightsquigarrow(j_{1},\dots,j_{k})$, one looks at the first coordinate where $i_{t}\neq j_{t}$ and then checks whether $i_{t}\rightsquigarrow j_{t}$ in the graph at the $t$th innermost level of the composition.
Hence, to study the limit as $k\to\infty$, we use a continuum digraph on the infinite product space $\Omega=[n]^{\mathbb{N}}$. Moreover, let $\rho$ be the infinite product of the uniform probability measure on $[n]$, which is a Radon measure on $\Omega$. Let $\mathcal{E}$ be the set of pairs $(\mathbf{i},\mathbf{j})\in\Omega\times\Omega$ such that if $t$ is the first index where $i_{t}\neq j_{t}$, then $i_{t}\rightsquigarrow j_{t}$ in $G$. Letting $E_{k}$ be the edge set of $G^{\circ k}$, we view $E_{k}\times\Omega$ as a subset of $\Omega$, where $E_{k}$ determines the values of the first $k$ coordinates. Then $E_{k+1}\times\Omega\subseteq E_{k}\times\Omega$, and we have
$\mathcal{E}=\bigcap_{k=1}^{\infty}E_{k}\times\Omega.$
Hence, by continuity of the measure,
$\lim_{k\to\infty}\rho(\mathcal{E}\Delta(E_{k}\times\Omega)).$
Therefore, by Proposition 1.2, we have the following result.
Theorem 5.9 (Limit theorem for iterated composition compare [10, Theorem 8.6], [9, Theorem 6.1]).
Fix $G\in\operatorname{Digraph}(n)$, and let $G^{\circ k}$ be its $k$-fold iterated composition. Let $\Omega=[n]^{\mathbb{N}}$ and let $\mathcal{E}$ be the edge set described above. Let $G^{\prime}$ be a rooted forest. Then
$\lim_{k\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G^{\circ k})|}{n^{k|V^{\prime}|}}=\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E})).$
Therefore, if $\mu_{k}\in\mathcal{P}(\mathbb{R})$ such that $\mu_{k}^{\uplus n^{k}}\to\mu$, then the limit $\widehat{\mu}=\lim_{k\to\infty}\boxplus_{G^{\circ k}}(\mu_{k})$ exists.
Remark 5.10.
Although this result is contained in [9, Theorem 6.1], the proof used here is different. In [9], the proof is based on showing the sequence of measures is Cauchy using the uniform continuity estimates for convolution, and there is no hope of generalizing this technique to the setting of Theorem 1.1. The proof of Theorem 1.1 here relies instead on the moment formulas as in [10], and thus also gives information about the moments of $\widehat{\mu}$ in the compactly supported case. Further, we remark that although the continuum limit construction and moment computations used here overlap with [10], we avoid the cumulant machinery of [10, §7].
5.4. Multi-regular digraphs
In [10], it was shown that for regular graphs the central limit distribution (under iterated composition) only depends on the number of vertices and the degree. For digraphs, “regular” means in this paper that the out-degree of each vertex is the same. Here we will generalize this result and consider sequences of multi-regular digraphs.
Fix $n\in\mathbb{N}$. Let $G_{n}=(V_{n},E_{n})$ be a digraph, and assume that
(5.3) $V_{n}=\bigsqcup_{j=1}^{m}V_{n,j},$
and further that for $i,j\in[n]$, there is a constant $A_{n,i,j}$ such that
(5.4) for $\text{for }v\in V_{n,i},\quad|\{w\in V_{n,j}:v\rightsquigarrow w\}|=A_{n,i,j},$
that is, each vertex in $V_{n,i}$ has $A_{n,i,j}$-many edges into $V_{n,j}$.
Let $G^{\prime}=(V^{\prime},E^{\prime})$ be a rooted tree. In order to compute $|\operatorname{Hom}(G^{\prime},G_{n})|$, we partition the set of homomorphisms based on which set $V_{n,j}$ contains the image of each vertex in $V^{\prime}$. More precisely, given a label function $\ell:V^{\prime}\to[m]$, let $\operatorname{Hom}_{\ell}(G^{\prime},G_{n})$ be the set of $\phi\in\operatorname{Hom}(G^{\prime},G_{n})$ such that $\phi(v)\in G_{n,\ell(v)}$ for all $v\in V^{\prime}$. Then $|\operatorname{Hom}_{\ell}(G^{\prime},G_{n})|$ can be computed by counting the number of choices for where to map each vertex of $G^{\prime}$ iteratively: For the root vertex $r$, there $|V_{n,\ell(r)}|$ choices for $\phi(r)$. For any non-root vertex $v$, let $v_{-}$ be its parent in the tree. Assuming that $\phi(v_{-})$ has already been chosen, then $v$ must be mapped to a vertex in $V_{n,\ell(v)}$ with an edge from $\phi(v_{-})\in V_{n,\ell(v_{-})}$, and hence there are $A_{n,\ell(v_{-}),\ell(v)}$ choices for $\phi(v)$. Therefore, we have
$|\operatorname{Hom}_{\ell}(G^{\prime},G_{n})|=|V_{n,\ell(r)}|\prod_{v\in V^{\prime}\setminus\{r\}}A_{n,\ell(v_{-}),\ell(v)}.$
(This argument can be formalized as an induction on $|V^{\prime}|$ where the inductive step considers removing one leaf from $G^{\prime}$.) Now summing over $\ell:V^{\prime}\to[n]$, we obtain
(5.5) $|\operatorname{Hom}(G^{\prime},G_{n})|=\sum_{\ell:V^{\prime}\to[n]}|V_{n,\ell(r)}|\prod_{v\in V^{\prime}\setminus\{r\}}A_{n,\ell(v_{-}),\ell(v)}.$
We remark that the output of this formula only depends on $|V_{n,j}|$’s and the $A_{n,i,j}$’s, and in particular any multi-regular digraph with these same constants will produce the same number of homomorphisms and hence the same convolution operation.
Next, we consider limits as $n\to\infty$. Assume that
(5.6) $\lim_{n\to\infty}\frac{|V_{n,j}|}{|V_{n}|}=t_{j}>0,\qquad\lim_{n\to\infty}\frac{A_{n,i,j}}{|V_{n}|}=a_{i,j}.$
Then we obtain from (5.5) that
$\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\sum_{\ell:V^{\prime}\to[n]}\frac{|V_{n,\ell(r)}|}{|V^{\prime}|}\prod_{v\in V^{\prime}\setminus\{r\}}\frac{A_{n,\ell(v_{-}),\ell(v)}}{|V^{\prime}|},$
hence we obtain the following result:
Proposition 5.11 (Limit theorem for multiregular digraphs).
Let $G_{n}=(V_{n},E_{n})$ be a multiregular graph satisfying (5.3) and (5.4) with respect to coefficients $A_{n,i,j}$, such that the limiting conditions (5.6) hold. Then
(5.7) $\lim_{n\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\sum_{\ell:V^{\prime}\to[m]}t_{\ell(r)}\prod_{v\in V^{\prime}\setminus\{r\}}a_{\ell(v_{-}),\ell(v)}=:\beta_{G^{\prime}}.$
Hence, by Theorem 1.1, if $\mu_{n}\in\mathcal{P}(\mathbb{R})$ and $\mu_{n}^{\uplus|V_{n}|}\to\mu$, then $\boxplus_{G_{n}}(\mu_{n})$ converges.
Next, let us describe how to compute the measure $\widehat{\mu}=\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})$ in this situation. We start with the fixed point equations in Proposition 4.4 for the $K$-transforms. Fix $n$. Then Proposition 4.4 gives a system of equations for $K_{\nu_{v}}$ where $\nu_{v}$ is the convolution with respect to $\operatorname{Walk}(G_{n},v)$ in the sense of [10]. Since the graph is multiregular, the isomorphism class of $\operatorname{Walk}(G_{n},v)$ is the same for all vertices $v$ in the same part $V_{n,j}$ of our partition. We denote by $\nu_{n,j}$ the common value of $\boxplus_{\operatorname{Walk}(G_{n},v)}(\mu)$ for $v\in V_{n,j}$. Then Proposition 4.4 yields
$K_{\nu_{n,i}}(z)$ $=K_{\mu_{n}}\left(z-\sum_{i=1}^{n}A_{n,i,j}K_{\nu_{n,j}}(z)\right)$
$K_{\boxplus_{G_{n}}(\mu_{n})}(z)$ $=\sum_{i=1}^{n}|V_{n,i}|K_{\nu_{n,i}}.$
Since we assume that $\mu_{n}^{\uplus 1/|V_{n}|}$ converges, we renormalize these equations as follows:
$|V_{n}|K_{\nu_{n,i}}(z)$ $=|V_{n}|K_{\mu_{n}}\left(z-\sum_{j=1}^{n}\frac{A_{n,i,j}}{|V_{n}|}\cdot|V_{n}|K_{\nu_{n,j}}(z)\right)$
$K_{\boxplus_{G_{n}}(\mu_{n})}(z)$ $=\sum_{j=1}^{n}\frac{|V_{n,i}|}{|V_{n}|}\cdot|V_{n}|K_{\nu_{n,i}}.$
By assumption, $|V_{n}|K_{\mu_{n}}\to K_{\mu}$. We will prove below that $|V_{n}|K_{\nu_{n,j}}(z)$ converges to some $K_{\nu_{j}}$, and that we can take the limit of the above equations.
Proposition 5.12 (Limit theorem for multiregular digraphs 2).
Let $G_{n}=(V_{n},E_{n})$ be a multiregular graph and assume (5.3), (5.4), (5.6). Let $\mu_{n}$ be a sequence of probability measures such that $\mu_{n}^{\uplus|V_{n}|}\to\mu$. Let $\widehat{\mu}=\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})$. Then there exist unique probability measures $(\nu_{i})_{i=1}^{m}$ satisfying
(5.8) $K_{\nu_{i}}(z)=K_{\mu}\left(z-\sum_{j=1}^{m}a_{i,j}K_{\nu_{j}}(z)\right),$
and in fact $\nu_{i}=\lim_{n\to\infty}\nu_{n,i}^{\uplus|V_{n}|}$ where $\nu_{i}$ is as above. Then the measure $\widehat{\mu}$ is given by
$K_{\widehat{\mu}}(z)=\sum_{i=1}^{m}t_{i}K_{\nu_{i}}(z).$
Proof.
The proof of convergence of $|V_{n}|K_{\nu_{n,i}}$ will follow roughly the same outline as the proof of Theorem 1.1. We first consider the case where $\mu_{n}=\mu^{\uplus 1/|V_{n}|}$ where $\mu$ is compactly supported, and then extend to the general case by equicontinuity.
First, recall from Proposition 4.4 that $\nu_{n,i}$ is the tree convolution of $\mu_{n}$ according to the tree $\operatorname{Walk}(G_{n},v)$ where $v$ is any vertex in $V_{n,i}$. Now let $\operatorname{Walk}(G_{n},V_{n,i})$ be the tree whose vertices are the (reverse) paths that start at some vertex in $V_{n,i}$. Note $\operatorname{Walk}(G_{n},V_{n,i})$ as the union of $\operatorname{Walk}(G_{n},v)$ for $v\in V_{n,i}$, where the root vertex $\emptyset$ is in their common intersection but otherwise they are disjoint. Thus,
$\boxplus_{\operatorname{Walk}(G_{n},V_{n,i})}(\mu_{n})=\biguplus_{v\in V_{n,i}}\boxplus_{\operatorname{Walk}(G_{n},v)}(\mu_{n})=\nu_{i}^{\uplus V_{n,i}}.$
Thus, in particular, letting $\tilde{\nu}_{n,i}=\boxplus_{\operatorname{Walk}(G_{n},V_{n,i})}(\mu_{n})$, we have $K_{\tilde{\nu}_{n,i}}=|V_{n,i}|K_{\nu_{n,i}}$. In terms of §3.1, $\operatorname{Walk}(G_{n},V_{n,i})$ produces a Hilbert space with summands $H_{v_{k}}^{\circ}\otimes\dots\otimes H_{v_{1}}^{\circ}$ where $v_{1}\in V_{n,i}$ rather than $v_{1}$ being arbitrary. Accordingly, the formula in Theorem 3.11 is changed to include only partitions $\pi$ where the outer blocks of $\pi$ are labeled by vertices in $V_{n,i}$ instead of arbitrary vertices. Hence, the formula for the moments of $\tilde{\nu}_{n,i}$ in this case is similar to Lemma 4.1 except that instead of all $\pi\in\mathcal{NC}_{k}(\ell,G_{n})$, we only take $\pi$ in the set $\mathcal{NC}_{k}(\ell,G_{n},V_{n,i})$ of all $\pi$ such that $\pi$ is compatible with $\ell$ and the labelling $\ell$ defines a homomorphism from $\mathrm{F}(\pi)\to G$ with all the outer blocks mapped to vertices in $V_{n,i}$. For a forest $G^{\prime}$, let us denote by $\operatorname{Hom}(G^{\prime},G_{n},V_{n,i})$ the set of homomorphisms from $G^{\prime}$ to $G_{n}$ such that all the root vertices of $G^{\prime}$ are mapped to vertices in $V_{n,i}$. Then as in Lemma 4.1, we get
$m_{k}(\tilde{\nu}_{n,i})$ $=\sum_{\pi\in\mathcal{NC}_{k}}|\operatorname{Hom}(\mathrm{F}(\pi),G_{n},V_{n,i})|\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu_{n}),$
$=\sum_{\pi\in\mathcal{NC}_{k}}\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G_{n},V_{n,i})|}{|V_{n}|^{|\pi|}}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu)$
since $\mu_{n}=\mu^{\uplus 1/|V_{n}|}$. Our counting argument for homomorphisms in the multiregular case implies that
$\frac{|\operatorname{Hom}(\mathrm{F}(\pi),G_{n},V_{n,i})|}{|V_{n}|^{|\pi|}}\to\sum_{\begin{subarray}{c}\ell:V^{\prime}\to[n]\\ \ell(r)=i\end{subarray}}t_{\ell(r)}\prod_{v\in V^{\prime}\setminus\{r\}}a_{\ell(v_{-}),\ell(v)},$
for the same reason as Proposition 5.11. Thus, by same reasoning as in Lemma 4.3 shows that $\tilde{\nu}_{n,i}$ converges as $n\to\infty$ to some $\tilde{\nu}_{i}$.
Now for the case of general $\mu_{n}$ such that $\mu_{n}^{\uplus|V_{n}|}\to\mu$, we use the fact that the mapping $\sigma\mapsto\boxplus_{\operatorname{Walk}(G_{n},V_{n,i})}(\sigma^{\uplus 1/|V_{n}|}$ is uniformly equicontinuous on any compact subset of $\mathcal{P}(\mathbb{R})$, which follows from [9, Theorem 6.2]. Thus, the same $3\epsilon$ argument from the proof of Theorem 1.1 in §4.2 applies here. Thus, we obtain convergence of $\tilde{\nu}_{n,i}$ in this case.
This means that $|V_{n,i}|K_{\nu_{n,i}}$ converges as $n\to\infty$ to $K_{\tilde{\nu}_{i}}$. Hence also $|V_{n}|K_{\nu_{n,i}}=\frac{|V_{n}|}{|V_{n,i}|}|V_{n,i}|K_{\nu_{n,i}}$ converges to $(1/t_{i})K_{\tilde{\nu}_{i}}$ as $n\to\infty$. Now let $\nu_{i}=\tilde{\nu}_{i}^{\uplus 1/t_{i}}$. Now recall that
$|V_{n}|K_{\nu_{n,i}}(z)=|V_{n}|K_{\mu_{n}}\left(z-\sum_{j=1}^{m}\frac{A_{n,i,j}}{|V_{n}|}\cdot|V_{n}|K_{\nu_{n,j}}(z)\right).$
We now know that $|V_{n}|K_{\nu_{n,i}}\to K_{\nu_{i}}$ as $n\to\infty$. Moreover, $|V_{n}|K_{\mu_{n}}$ converges to $K_{\mu}$. Also, the functions $|V_{n}|K_{\mu_{n}}$ are equicontinuous because the measures $\mu_{n}^{\uplus|V_{n}|}$ inhabit a precompact subset of $\mathcal{P}(\mathbb{R})$ since $\mu_{n}^{\uplus|V_{n}|}\to\mu$. These facts together imply by a $3\epsilon$ argument that we can take the limit of the above equation and obtain
$K_{\nu_{i}}(z)=K_{\mu}\left(z-\sum_{j=1}^{m}a_{i,j}K_{\nu_{j}}(z)\right).$
Similar reasoning shows that we can take the limit of the equation $K_{\boxplus_{G_{n}}(\mu_{n})}=\sum_{i=1}^{m}|V_{n,i}|K_{\nu_{n,i}}$ to obtain $K_{\widehat{\mu}}=\sum_{i=1}^{m}t_{i}K_{\nu_{i}}$. ∎
Proposition 5.12 allows for numerically tractable computations of limit measures associated to multiregular digraphs $G_{n}$.
Example 5.13 (Central limit distribution for multiregular digraphs).
Suppose we want to find the central limit distribution. Since the Boolean central limit distribution is $\mu=(1/2)(\delta_{-1}+\delta_{1})$, the central limit distribution for the sequence $G_{n}$ will be given by the corresponding measure $\widehat{\mu}$. Note that in this case $K_{\mu}(z)=1/z$. Thus, (5.8) reduces to
$K_{\nu_{i}}(z)=\left(z-\sum_{j=1}^{n}a_{i,j}K_{\nu_{j}}(z)\right)^{-1},$
or equivalently
$zK_{\nu_{i}}(z)-\sum_{j=1}^{n}a_{i,j}K_{\nu_{i}}(z)K_{\nu_{j}}(z)=1.$
In other words, the $m$ unknowns $(K_{\nu_{1}}(z),\dots,K_{\nu_{m}}(z))$ satisfy a quadratic system of $m$ equations. Then $\widehat{\mu}$ is obtained by $K_{\widehat{\mu}}=\sum_{i=1}^{n}t_{i}K_{\nu_{i}}$.
Figures 1, 2, 3 show examples of numerical approximations of central limit densities using the fixed point equation (5.8). Changing the parameters produces symmetric distributions whose shape can be semicircular, become more flat, and then develop a concave shape in the middle with two bumps on the boundary, somewhat resembling the arcsine distribution. Of course, for special cases of a complete graph and its complement, one obtains the semicircular distribution and the Bernoulli distribution respectively.
1.51.5-1.5- 1.51.51.51.51.50.40.40.40.4
Figure 1. Approximation of the central limit density for $2$-regular digraphs with $t_{1}=0.3$, $t_{2}=0.7$, $a_{1,1}=0.2$, $a_{1,2}=0.4$, $a_{2,1}=0.2$, $a_{2,2}=0.5$. We approximated the density using the imaginary part of the Cauchy transform at $x+iy$ where $y=0.001$, and we approximated the Cauchy transform using $10,000$ iterations of the fixed point equation.
1.51.5-1.5- 1.51.51.51.51.51.01.01.01.0
Figure 2. Approximation of the central limit density for $2$-regular digraphs with $t_{1}=0.3$, $t_{2}=0.7$, $a_{1,1}=0.2$, $a_{1,2}=0.4$, $a_{2,1}=0.2$, $a_{2,2}=0.1$. We use $y=0.0001$ and $50,000$ iterations.
22-2- 222220.40.40.40.4
Figure 3. Approximation of the central limit density for $2$-regular digraphs with $t_{1}=0.5$, $t_{2}=0.5$, $a_{1,1}=0.4$, $a_{1,2}=0.5$, $a_{2,1}=0.4$, $a_{2,2}=0.5$. We used $y=0.0001$ and $50,000$ iterations.
Example 5.14 (Poisson distribution for multiregular digraphs).
Now consider the Poisson limit theorem or law of small numbers. Recall from Example 5.8 that the boolean analog of the Poisson distribution of intensity $\lambda$ is $\mu=\frac{1}{1+\lambda}\delta_{0}+\frac{\lambda}{1+\lambda}\delta_{1+\lambda}$, and its $K$-transform is $K_{\mu}(z)=\lambda z/(z-1)$. Thus, (5.8) becomes
$K_{\nu_{i}}(z)=\lambda\left(z-\sum_{j=1}^{n}a_{i,j}K_{\nu_{j}}(z)\right)\left(z-\sum_{j=1}^{n}a_{i,j}K_{\nu_{j}}(z)-1\right)^{-1},$
which after algebraic manipulation can be written equivalently as
$K_{\nu_{i}}(z)=(K_{\nu_{i}}(z)-\lambda)\left(z-\sum_{j=1}^{n}a_{i,j}K_{\nu_{j}}(z)\right).$
Thus, similar to the central limit case, $K_{\nu_{i}}(z)$’s satisfy a quadratic system of $m$ equations in $m$ unknowns.
Figures 4 and 5 show numerical approximations of Poisson limit distributions using (5.8).
22224444
Figure 4. Approximation of the Poisson limit density with $\lambda=1$ for $2$-regular digraphs. We computed using the same parameters as in Figure 1, except on the interval $[0,0.1]$, we used a $y=0.00001$ and $100,000$ iterations. The computation at $x=0$ suggests that the measure to have an atom at 00 of mass about $0.25$.
33336666
Figure 5. Approximation of the Poisson limit density with $\lambda=3$ for $2$-regular digraphs. We computed using the same parameters as in Figure 1.
Example 5.15 (Cauchy distribution for multiregular digraphs).
If we take $\mu$ to be the standard Cauchy distribution, then $K_{\mu}(z)=-i$ in the upper half plane. Thus, (5.8) tells us that $K_{\nu_{i}}(z)=-i$. Hence, $K_{\widehat{\mu}}(z)=\sum_{i=1}^{m}t_{i}K_{\nu_{i}}(z)=-i$ since $t_{1}+\dots+t_{m}=1$. It follows that when $\mu$ is the standard Cauchy distribution, then also $\widehat{\mu}$ is the standard Cauchy distribution.
5.5. Sparse graphs
The next proposition shows that if a sequence of digraphs is sufficiently sparse, then the normalized count of homomorphisms converges to $\beta_{G^{\prime}}=0$, and so the $G_{n}$-convolution is asymptotically boolean convolution. This is a generalization of the case of bm-independence for posets given by regular trees from [28, §8].
Proposition 5.16 (Limit theorem for sparse graphs).
Let $G_{n}$ be a sequence of digraphs such that $|E_{n}|/|V_{n}|^{2}\to 0$. Then for every rooted tree $G^{\prime}$ with more than one vertex, we have
$\lim_{n\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=0.$
Hence, in this case, if $\mu_{n}^{\uplus|V_{n}|}\to\mu$, then $\boxplus_{G_{n}}(\mu_{n})\to\mu$ also.
Proof.
Suppose that $G^{\prime}$ is a rooted tree with more than one vertex. Let $r$ be the root, and fix some $v$ which is a child of the root. Given a homomorphism $\phi:G^{\prime}\to G_{n}$, the pair $(r,v)$ must be mapped to some edge. There are $|E_{n}|$ choices for this edge. Then the remaining $|V^{\prime}|-2$ vertices must be mapped to some vertex in $V_{n}$, and so the number of choices for the rest of the values of $\phi$ is at most $|V^{\prime}|^{|V_{n}|-2}$. Thus,
$|\operatorname{Hom}(G^{\prime},G_{n})|\leq|E_{n}||V^{\prime}|^{|V_{n}|-2},$
so
$\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}\leq\frac{|E_{n}|}{|V_{n}|^{2}}\to 0.$
Hence, $\beta_{G^{\prime}}=0$ unless $G^{\prime}$ has only one vertex. More generally, if $G^{\prime}$ is a rooted forest, then $\beta_{G^{\prime}}=0$ unless $G^{\prime}$ has no edges, in which case $\beta_{G^{\prime}}=1$.
It follows that in Lemma 4.3, the moments $\boxplus_{G}(\mu^{\uplus 1/|V_{n}|})$ in (4.1) are given by the sum over interval partitions of $\kappa_{\operatorname{bool},\pi}(\mu)$, since $\mathrm{F}(\pi)$ has no edges if and only if $\pi$ is interval partitions. This means that when we take $\mu_{n}=\mu^{\uplus 1/|V_{n}|}$, then the moments of the limiting measure in (4.1) are the same as the moments of $\mu$.
The general statement that if $\mu_{n}^{\uplus|V_{n}|}\to\mu$, then $\boxplus_{G_{n}}(\mu_{n})\to\mu$ follows from the equicontinuity of the convolution operations as in §4.2. ∎
Example 5.17 (Posets given by finite trees).
The following example is from [28, §8]. Fix $d$. Let $T_{n}$ be the $d$-regular rooted tree (where each vertex has $d$ children) truncated to depth $n$. Let $G_{n}=(V_{n},E_{n})$ be the graph where $v\rightsquigarrow w$ if $v$ is an ancestor of $w$ in the tree $T_{n}$ (i.e. the edge relation for $G_{n}$ is the transitive closure of the edge relation for $T_{n}$). Observe that
$|V_{n}|=\sum_{j=0}^{n}d^{j}=\frac{d^{n+1}-1}{d-1}.$
Meanwhile, the number of edges can be counted as follows: Each vertex at depth $j$ in the tree has $\sum_{i=1}^{n-j}d^{i}$ descendants, which evaluates to $d(d^{n-j}-1)/(d-1)$. Now summing this over all the vertices, we obtain
$|E_{n}|=\sum_{j=0}^{n}d^{j}\cdot\frac{d(d^{n-j}-1)}{d-1}=\frac{d}{d-1}\sum_{j=0}^{n}(d^{n}-d^{j})\leq\frac{n(d^{n+1}-1)}{d-1}=n|V_{n}|.$
Thus,
$\frac{|E_{n}|}{|V_{n}|^{2}}\leq\frac{n}{|V_{n}|}=\frac{n(d-1)}{(d^{n+1}-1)}\to 0.$
Hence, the limiting measures for the sequence $G_{n}$ reduces to those of the boolean case. This generalizes the observation of [28] that the central limit measure for this case is $(1/2)(\delta_{-1}+\delta_{1})$.
Remark 5.18.
Our argument to bound the number of homomorphisms in the proof of Proposition 5.16 generalizes to yield the following statement: If $G^{\prime}$ is a rooted tree, $G^{\prime\prime}$ is a rooted subtree of it, and $G$ is any finite digraph, then
$\frac{|\operatorname{Hom}(G^{\prime},G)|}{|V|^{|V^{\prime}|}}\leq\frac{|\operatorname{Hom}(G^{\prime\prime},G)|}{|V|^{|V^{\prime\prime}|}}.$
The reason for this is that every homomorphism $G^{\prime}\to G$ restricts to a homomorphism $G^{\prime\prime}\to G$. Hence,
$\operatorname{Hom}(G^{\prime},G)\subseteq\operatorname{Hom}(G^{\prime\prime},G)\times V^{V^{\prime}\setminus V^{\prime\prime}},$
where $V^{V^{\prime}\setminus V^{\prime\prime}}$ denotes the set of all functions $V^{\prime}\setminus V^{\prime\prime}\to V$. Thus, we get
$|\operatorname{Hom}(G^{\prime},G)|\leq|\operatorname{Hom}(G^{\prime\prime},G)||V|^{|V^{\prime}|-|V^{\prime\prime}|},$
which is the inequality asserted above. In particular, in the situation where $\beta_{G^{\prime}}=\lim_{n\to\infty}|\operatorname{Hom}(G^{\prime},G_{n})|/|V_{n}|^{|V^{\prime}|}$ exists for all rooted trees $G^{\prime}$, then we have $\beta_{G^{\prime}}\leq\beta_{G^{\prime\prime}}$ whenever $G^{\prime\prime}$ is a rooted subtree of $G^{\prime}$ (in fact, we do not even need the root of $G^{\prime\prime}$ to agree with the root of $G^{\prime}$).
6. Fock space models
Proposition 1.2 described a general situation when the limit $\beta_{G^{\prime}}$ exist based on discrete approximations of measurable digraphs. Based on this construction, we want to describe the limiting measures in Theorem 1.1 (in the compactly supported case) as the spectral distributions of certain operators on a Fock space, which is a continuum analog of the $G$-free product space studied in §3. We remark that a similar idea was explored in the free case in [2]. Our Fock space is a direct sum of terms $L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k})$ for some measurable space $\Omega$, measure $\rho_{k}$ on $\Omega^{\times k}$, and Hilbert space $\mathcal{H}$. The operators will have the form $\mathfrak{n}(\phi)+\ell(h)+\ell^{*}(h)+\mathfrak{m}(S)$, where $\ell(h)$ and $\ell(h)^{*}$ are creation and annihilation operators associated to some $h\in L^{2}(\Omega,\rho_{1};\mathcal{H})$, and $\mathfrak{n}(\phi)$ is a multiplication operator associated to $\phi\in L^{\infty}(\Omega)$, and $\mathfrak{m}(S)$ is another type of multiplication operator associated to $S\in L^{\infty}(\Omega;B(\mathcal{H}))$.
6.1. Construction of a Fock space and operators thereon
Although in the last section, we considered $(\Omega,\rho)$ to be a probability measure space, here we will proceed more generally with a complete $\sigma$-finite measure space, in order to include such examples as the Fock spaces supporting Brownian motions on $[0,\infty)$. Moreover, while in §5, we considered edges given by $\mathcal{E}\subseteq\Omega\times\Omega$, we now introduce a weighted version where a general nonnegative $w:\Omega\times\Omega\to[0,\infty)$ replaces the indicator function $\mathbf{1}_{\mathcal{E}}$.
Construction 6.1.
Let $(\Omega,\rho)$ be a complete $\sigma$-finite measure space, and recall that there is a unique complete measure space $\Omega^{\times k},\rho^{\times k}$ obtained from the product measure construction. Let $w\in L^{\infty}(\Omega\times\Omega)$ with $w\geq 0$. For $k\geq 1$, let $\rho_{k}$ be the measure on $\Omega^{\times k}$ given by
$d\rho_{k}(\omega_{1},\dots,\omega_{k})=w(\omega_{k},\omega_{k-1})\dots w(\omega_{2},\omega_{1})\,d\rho^{\times k}(\omega_{1},\dots,\omega_{k});$
here $\rho_{1}=\rho$. Note that $(\Omega^{\times k},\rho_{k})$ can be completed to a complete measure space. Let $\mathcal{H}$ be a Hilbert space. Then we define the Fock space as the Hilbert space
$\mathcal{F}(\Omega,\rho,w,\mathcal{H})=\mathbb{C}\oplus\bigoplus_{k\in\mathbb{N}}L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k}).$
We denote the vector $1$ in the first summand by $\xi$. Furthermore, we adopt the convention that $\Omega^{\times 0}$ is a single point, $\rho_{0}$ is the unique probability measure on it, and $\mathcal{H}^{\otimes 0}=\mathbb{C}$; thus, $L^{2}(\Omega^{\times 0},\rho_{0},\mathcal{H}^{\times 0})=\mathbb{C}$.
Construction 6.2.
Consider the same setup as in the previous construction. Let $h\in L^{2}(\Omega,\rho;\mathcal{H})$. Then we define the left creation operator $\ell(h):\mathcal{F}(\Omega,\rho,w,\mathcal{H})\to\mathcal{F}(\Omega,\rho,w,\mathcal{H})$ as follows. For $k\geq 0$ and $f\in L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k})$, let
$[\ell(h)f](\omega_{1},\dots,\omega_{k+1})=h(\omega_{1})\otimes f(\omega_{2},\dots,\omega_{k+1}).$
This formula immediately yields a well-defined element of $L^{2}(\Omega^{\times(k+1)},\rho_{1}\times\rho_{k};\mathcal{H}^{\otimes(k+1)})$, but in fact it even yields a well-defined element of $L^{2}(\Omega^{\times(k+1)},\rho_{k+1};\mathcal{H}^{\otimes(k+1)})$ since
[…]
and moreover
$\lVert\ell(h)f\rVert_{L^{2}(\Omega^{\times(k+1)},\rho_{k+1},\mathcal{H})}\leq\lVert w\rVert_{L^{\infty}(\Omega\times\Omega)}\lVert h\rVert_{L^{2}(\Omega,\rho;\mathcal{H})}\lVert f\rVert_{L^{2}(\Omega^{\times k},\rho_{k})}.$
(Note that in the $n=0$ case, we have $\ell(h)\xi=h\in L^{2}(\Omega,\rho;\mathcal{H})$.) It follows that $\ell(h)$ defines a bounded operator on $\mathcal{F}(\Omega,\rho,w,\mathcal{H})$ with
$\lVert\ell(h)\rVert\leq\lVert w\rVert_{L^{\infty}(\Omega\times\Omega)}\lVert h\rVert_{L^{2}(\Omega,\rho;\mathcal{H})}.$
Thus, the creation operator is well-defined. Its adjoint $\ell(h)^{*}$ is called the left annihilation operator associated to $h$.
Observation 6.3.
The annihilation operator $\ell(h)$ satisfies
$\ell(h)^{*}\xi=0$
Moreover, $\ell(h)^{*}$ maps $L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k})$ into $L^{2}(\Omega^{\times(k-1)},\rho_{k-1};\mathcal{H}^{\otimes(k-1)})$ for each $k\geq 1$ and satisfies
[…]
where $\langle h(\omega),-\rangle\otimes\operatorname{id}_{\mathcal{H}^{\otimes(k-1)}}$ denotes the map
$\mathcal{H}^{\otimes k}\to\mathcal{H}^{\otimes(k-1)}:f_{1}\otimes\dots\otimes f_{k}\mapsto\langle h(\omega),f_{1}\rangle_{\mathcal{H}}f_{2}\otimes f_{3}\otimes\dots\otimes f_{k},$
and in the case $k=1$, it is $f\mapsto\langle h(\omega),f\rangle_{\mathcal{H}}\in\mathbb{C}$.
Remark 6.4.
It may be easier to understand the annihilation operator through its action on simple tensors. If $f(\omega_{1},\dots,\omega_{k})=f_{1}(\omega_{1})\otimes\dots\otimes f_{k}(\omega_{k})$, where $f_{j}\in L^{2}(\Omega,\rho)$, then
[…]
Definition 6.5.
$L^{\infty}(\Omega,B(L^{2}(\Omega,\rho))$ denotes the space of essentially bounded *-SOT measurable maps from $\Omega$ into $B(L^{2}(\Omega,\rho))$.
Construction 6.6.
Consider the Fock space defined above. Let $S\in L^{\infty}(\Omega,B(\mathcal{H}))$. Then we define the multiplication operator $\mathfrak{m}(S):\mathcal{F}(\Omega,\rho,w,\mathcal{H})\to\mathcal{F}(\Omega,\rho,w,\mathcal{H})$ by $\mathfrak{m}(T)|_{\mathbb{C}}=0$ and for $k\geq 1$ and $f\in L^{2}(\Omega^{\times k},\rho_{k})$,
$[\mathfrak{m}(S)f](\omega_{1},\dots,\omega_{k}):=(S(\omega_{1})\otimes\operatorname{id}_{\mathcal{H}^{\otimes(n-1)}})(f(\omega_{1},\dots,\omega_{k})).$
Since $S(\omega_{1})\otimes\operatorname{id}_{\mathcal{H}^{\otimes(k-1)}}$ defines a bounded operator on $\mathcal{H}^{\otimes k}$ with norm less than or equal to that of $S(\omega_{1})$, we deduce that
$\lVert\mathfrak{m}(S)f\rVert_{L^{2}(\Omega^{\times k},\rho_{k})}\leq\lVert S\rVert_{L^{\infty}(\Omega,B(L^{2}(\Omega,\rho))}\lVert f\rVert_{L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k})}.$
Thus, $\mathfrak{m}(S)$ defines a bounded operator on $L^{2}(\Omega^{\times k},\rho_{k})$ with norm less than or equal to that of $\lVert S\rVert_{L^{\infty}(\Omega,B(\mathcal{H}))}$. Since $\mathcal{F}(\Omega,\rho,w,\mathcal{H})$ is the direct sum of the subspaces $L^{2}(\Omega^{\times n},\rho_{k};\mathcal{H}^{\otimes k})$, we conclude that $\mathfrak{m}(S)$ is a bounded operator on the Fock space.
Observation 6.7.
$\mathfrak{m}:L^{\infty}(\Omega,B(\mathcal{H}))\to B(\mathcal{F}(\Omega,\rho,w,\mathcal{H}))$ is a *-homomorphism.
Construction 6.8.
Let $\phi\in L^{1}(\Omega,\rho)$. Then we define an operator $\mathfrak{n}(\phi)$ on $\mathcal{F}(\Omega,\rho,w\mathcal{H})$ by
[…]
and for $f\in L^{2}(\Omega^{\times k},\rho_{k},\mathcal{H}^{\otimes k})$,
[…]
Here $\mathfrak{n}(\phi)$ maps $L^{2}(\Omega^{\times k},\rho_{k},\mathcal{H}^{\otimes k})$ into itself for each $k$. Also, $\lVert\mathfrak{n}(\phi)\rVert\leq\lVert w\rVert_{L^{\infty}(\Omega^{\times 2},\rho^{\times 2})}\lVert\phi\rVert_{L^{1}(\Omega,\rho)}$.
The following observation may be helpful for understanding the motivation or intuition of the operator $\mathfrak{n}(\phi)$.
Observation 6.9.
Let $h_{1},h_{2}\in\mathcal{H}$, let $\psi_{1},\psi_{2}\in L^{2}(\Omega,\rho)$, and let $\psi_{j}h_{j}\in L^{2}(\Omega,\rho;\mathcal{H})$ be the map $\omega\mapsto\psi_{j}(\omega)h_{j}$. Then
$\ell(\psi_{1}h_{1})^{*}\ell(\psi_{2}h_{2})=\langle h_{1},h_{2}\rangle_{\mathcal{H}}\mathfrak{n}(\overline{\psi}_{1}\psi_{2}).$
This is proved by directly computing the effect of these operators on some $f\in L^{2}(\Omega^{\times n},\rho_{k};\mathcal{H}^{\otimes k})$.
6.2. Combinatorial formula for operators on a Fock space
Our next goal is derive a combinatorial formula for the “joint moment”
$\langle\xi,T_{n}T_{n-1}\dots T_{1}\xi\rangle,$
where $T_{1}$, …, $T_{n}$ are creation, annihilation, or multiplication operators on the Fock space $\mathcal{F}(\Omega,\rho,w,\mathcal{H})$. For the sake of induction, we will find a combinatorial expression for the vector $T_{n}T_{n-1}\dots T_{1}\xi$ itself.
Setup: Let $T_{1}$, …, $T_{n}$ be operators on $\mathcal{F}(\Omega,\rho,w,\mathcal{H})$ such that each $T_{j}$ is one of the following types:
  • $T_{j}=\ell(h_{j})$ for some $h_{j}\in L^{2}(\Omega,\rho,\mathcal{H})$,
  • $T_{j}=\ell(h_{j})^{*}$ for some $h_{j}\in L^{2}(\Omega,\rho,\mathcal{H})$.
  • $T_{j}=\mathfrak{m}(S_{j})$ for some $S_{j}\in L^{\infty}(\Omega,\rho;B(\mathcal{H}))$.
  • $T_{j}=\mathfrak{n}(\phi_{j})$ for some $\phi_{j}\in L^{1}(\Omega,\rho)$.
Let $k(j)$ be the number of creation operators among $\{T_{1},\dots,T_{j}\}$ minus the number of annihilation operators among $\{T_{1},\dots,T_{j}\}$.
Observation 6.10.
Because a creation operator maps $L^{2}(\Omega^{\times k},\rho_{k};\mathcal{H}^{\otimes k})$ into $L^{2}(\Omega^{\times(k+1)},\rho_{k+1};\mathcal{H}^{\otimes(k+1)})$ while an annihilation operator does the opposite, one can verify by induction on $n$ that $T_{n}\dots T_{1}\xi\in L^{2}(\Omega^{\times k(n)},\rho_{k(n)};\mathcal{H}^{\otimes k(n)})$ if $k(j)\geq 0$ for all $j$. Moreover, $T_{n}\dots T_{1}\xi=0$ if $k(j)$ is ever negative.
Now assume that $k(j)\geq 0$ for all $j$. For each $j=1$, …, $n$, we define $m(j)$ as follows:
  • If $T_{j}=\mathfrak{n}(\phi_{j})$ or $T_{j}=\ell(h_{j})$, then set $m(j)=j$.
  • If $T_{j}=\ell(h_{j})^{*}$ or $T_{j}=\mathfrak{m}(S_{j})$, let $m(j)$ be the greatest index $m$ such that $k(m-1)<k(j-1)$.
In the second case, note that $k(j-1)\geq k(j)$. By definition, $k(i)\geq k(j-1)$ for all $i$ between $m(j)$ and $j$. Moreover, since $|k(i+1)-k(i)|\leq 1$, we deduce that $k(m-1)=k(m)-1=k(j-1)-1$, and thus $T_{m(j)}$ is a creation operator.
Remark 6.11.
The intuition behind the choice of $m(j)$ is the following: Each creation operator “creates” a new particle that is tensored onto the left of the vector it acts on, while each annihilation operator “annihilates” a particle. The multiplication operators neither create nor annihilate anything. The number $k(j)$ represents the current tensor degree, or the number of particles that exist at time $j$ (after the application of $T_{j}$). If $T_{j}$ is $\ell(h_{j})^{*}$ or $\mathfrak{m}(S_{j})$, then $T_{m(j)}$ is the creation operator that created the newest particle that still exists, the one that $T_{j}$ is acting on. In the case where $T_{j}=\mathfrak{n}(\phi_{j})$, in light of Observation 6.9, we can imagine that $T_{j}$ creates and immediately annihilates some ephemeral particle. With respect to this picture, the next construction will be to group together all the operators that act on “the same particle”; the indices of these operators will form the block of a non-crossing partition — non-crossing because the operators can only act on the newest existing particle, and thus this particle must be annihilated before any operator can act on an older particle.
Let $\pi$ be the partition of $[n]$ given by $i\sim_{\pi}j$ if and only if $m(i)=m(j)$. Observe that
  • Each creation operator satisfies $m(j)=j$ and hence is the first element of its block.
  • If $T_{j}$ is an annihilation operator, then it is the last element of its block. This is because $k(j)<k(j-1)$ and this prevents any later index $i>j$ from having $m(i)=m(j)$.
  • If $T_{j}=\mathfrak{n}(\phi_{j})$, then $\{j\}$ is a singleton block of $\pi$.
We call a block finished if either has a single $\mathfrak{n}$ operator or has both a creation and annihilation operator. Otherwise, we call a block unfinished. The unfinished blocks will have a creation operator but no annihilation operator.
Observation 6.12.
For every sequence of creation, annihilation, and multiplication operators with $k(j)\geq 0$ for all $j$, the associated partition $\pi$ is non-crossing.
Proof.
Suppose for contradiction that $i<j<i^{\prime}<j^{\prime}$ with $i\sim_{\pi}i^{\prime}$ and $j\sim_{\pi}j^{\prime}$ for $i\not\sim_{\pi}j$. By the preceding discussion, since $i\sim i^{\prime}$ and $i<j<i^{\prime}$, we must have $k(j)\geq k(i)=k(i^{\prime}-1)$. Note that $T_{j^{\prime}}$ must be an annihilation operator or $\mathfrak{m}$ operator, and $k(j^{\prime}-1)=k(j)$. If $k(j)>k(i)$, then $k(i^{\prime}-1)<k(j^{\prime}-1)$, which would imply that $m(j^{\prime})\geq i^{\prime}$ by definition of $m$, but this contradicts the fact that $m(j^{\prime})=m(j)\leq j<i^{\prime}$. On the other hand, suppose $k(j)=k(i)=k(i^{\prime}-1)$. Since $j\sim_{\pi}j^{\prime}$, $T_{j}$ cannot be an annihilation operator, so $k(j-1)\leq k(j)$, but also $k(j-1)\geq k(i)=k(j)$, since $j-1$ is between $i$ and $i^{\prime}$, hence $k(j-1)=k(j)$. Recall that $m(i^{\prime})=m(i)\leq i$ is the last index before $i^{\prime}$ where $k(m-1)<k(i^{\prime}-1)=k(j)$. Since $k(t)\geq k(m(i))$ for all $t$ between $m(i)=m(i^{\prime})$ and $i^{\prime}$, we deduce that $m(i^{\prime})$ is also the last index before $j$ where $k(m-1)<k(j-1)=k(j)$, which implies that $m(j)=m(i)$, which contradicts the assumption that $i$ and $j$ are in different blocks of $\pi$. ∎
With the notation above, for each unfinished block $B=\{i_{1},\dots,i_{|B|}\}$, let $h_{B}\in L^{2}(\Omega,\rho)$ be given by
$h_{B}(\omega)=S_{i_{|B|}}(\omega)\dots S_{i_{2}}(\omega)h_{i_{1}}(\omega).$
For each finished block $B=\{i_{1},\dots,i_{|B|}\}$, let
$\phi_{B}(\omega)=\begin{cases}\phi_{i_{1}},&|B|=1,\\ \langle h_{i_{|B|}}(\omega),S_{i_{|B|-1}}(\omega)\dots S_{i_{2}}(\omega)h_{i_{1}}(\omega)\rangle_{\mathcal{H}}.\end{cases}$
Proposition 6.13.
Let $T_{n}\dots T_{1}$ be a sequence of creation, annihilation, $\mathfrak{m}$, and $\mathfrak{n}$ operators as above, such that $k(j)\geq 0$ for all $j$. Let $\operatorname{pred}(B)$ denote the predecessor of $B$ in $\pi$. Let $B_{1}$, …, $B_{s}$ be the unfinished blocks of $\pi$ listed so that $\min B_{1}<\dots<\min B_{t}$. Then $t=k(n)$ and $T_{n}T_{n-1}\dots T_{1}\xi\in L^{2}(\Omega^{\times t},\rho_{t};\mathcal{H}^{\otimes t})$ and
(6.1) $[T_{n}T_{n-1}\dots T_{1}\xi](\omega_{B_{t}},\dots,\omega_{B_{1}})\\ =\int_{\Omega^{\times\{B\text{ finished}\}}}\prod_{\begin{subarray}{c}B\text{ finished}\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{\operatorname{pred}(B)},\omega_{B})\prod_{B\text{ finished}}[\phi_{B}(\omega_{B})\,d\rho(\omega_{B})]\\ h_{B_{t}}(\omega_{B_{t}})\otimes\dots\otimes h_{B_{1}}(\omega_{B_{1}})$
In the case $t=0$, we interpret $h_{B_{t}}(\omega_{B_{t}})\otimes\dots\otimes h_{B_{1}}(\omega_{B_{1}})$ as $\xi$.
Proof.
We proceed by induction on $n$. The base case is $n=0$, for which both sides reduce to $\xi$.
For the induction step, suppose the claim holds for $T_{n}T_{n-1}\dots T_{1}$, and we will prove it for $T_{n+1}T_{n}\dots T_{1}$. Let $\pi^{\prime}$ be the partition associated to $T_{n+1}$, …, $T_{1}$. By restricting $\pi^{\prime}$ to $\{n,\dots,1\}$, we obtain the partition $\pi$ associated to $T_{n}$, …, $T_{1}$.
  • If $T_{n+1}=\ell(h_{n+1})$, then $m(n+1)=n+1$ and $\{n+1\}$ is an unfinished block in $\pi^{\prime}$, so $\pi^{\prime}=\pi\cup\{\{n+1\}\}$. If $B_{1}$, …, $B_{t}$ are the unfinished blocks in $\pi$, then the unfinished blocks of $\pi^{\prime}$ will be $B_{1}$, …, $B_{t}$ and $B_{t+1}:=\{n+1\}$. Moreover, $h_{B_{t+1}}=h_{n+1}$. Therefore, the right-hand side o (6.1) for $\pi^{\prime}$ will be the same as the right-hand side of (6.1) for $\pi$ except with $h_{B_{t+1}}(\omega_{B_{t+1}})$ tensored onto the front of $h_{B_{t}}(\omega_{B_{t}})\otimes\dots\otimes h_{B_{1}}(\omega_{B_{1}})$. Meanwhile,
    $T_{n+1}[T_{n}\dots T_{1}\xi](\omega_{B_{t+1}},\dots,\omega_{B_{1}})=h_{n+1}(\omega_{B_{t+1}})\otimes[T_{n}\dots T_{1}](\omega_{B_{t}},\dots,\omega_{B_{1}}),$
    and hence (6.1) will be true for $T_{n+1}$, …, $T_{1}$.
  • Suppose that $T_{n+1}=\ell(h_{n+1})^{*}$. Then $n+1$ will be the last element of a finished block $B^{\prime}$ in $\pi^{\prime}$ and $B=B^{\prime}\setminus\{n+1\}$ will be an unfinished block in $\pi$. Write $B=\{i_{1},\dots,i_{s}\}$. Then
    $\phi_{B^{\prime}}(\omega_{B^{\prime}})=\langle h_{n+1}(\omega_{B^{\prime}}),S_{i_{s}}(\omega_{B^{\prime}})\dots S_{i_{2}}(\omega_{B^{\prime}})h_{i_{1}}(\omega_{B^{\prime}})\rangle=\langle h_{n+1}(\omega_{B^{\prime}},h_{B}(\omega_{B^{\prime}})\rangle.$
    where $h_{B}$ is the vector corresponding to $B$ as an unfinished block of $\pi$. Thus, to obtain the right-hand side of (6.1) for $\pi^{\prime}$ from the right-hand side of (6.1) for $\pi$, one removes the term $h_{B}(\omega_{B})$ from the unfinished blocks and adds the term $\phi_{B^{\prime}}(\omega_{B^{\prime}})\,d\rho(\omega_{B^{\prime}})$ to the finished blocks along with $w(\omega_{B^{\prime}},\omega_{\operatorname{pred}(B^{\prime})})$ if $\operatorname{depth}(B^{\prime})>1$. Meanwhile, looking at the left-hande side of (6.1) the application of $\ell(h_{n+1})$ to $T_{n}\dots T_{1}\xi$ will precisely pair $h_{n+1}(\omega_{B^{\prime}})$ with $h_{B}(\omega_{B^{\prime}})$ in $\mathcal{H}$, multiply by $w(\omega_{\operatorname{pred}(B^{\prime})},\omega_{B^{\prime}})$ if $\operatorname{depth}(B^{\prime})>1$, and then integrate $d\rho(\omega_{B^{\prime}})$. Hence, the left- and right-hand sides of (6.1) agree for $\pi^{\prime}$.
  • Suppose that $T_{n+1}=\mathfrak{n}(\phi_{n+1})$. Then $B^{\prime}:=\{n+1\}$ is a new finished block in $\pi^{\prime}$ and $\phi_{B^{\prime}}=\phi_{n+1}$. The right-hand side of (6.1) for $\pi^{\prime}$ differs from that for $\pi$ by adding a new term $\phi_{B^{\prime}}(\omega_{B^{\prime}})d\rho(\omega_{B^{\prime}})$ to the finished blocks, along with $w(\omega_{\operatorname{pred}(B^{\prime})},\omega_{B^{\prime}})$ if $\operatorname{depth}(B^{\prime})>1$. This agrees with what happens when we apply the operator $\ell(h_{n+1})^{*}$ to $T_{n}\dots T_{1}\xi$.
  • Finally, suppose that $T_{n+1}=\mathfrak{m}(S_{n+1})$. Then $n+1$ is an element of some unfinished block $B^{\prime}$ of $\pi^{\prime}$ such that $B=B^{\prime}\setminus\{n+1\}$ is also unfinished in $\pi$. Write $B=\{i_{1},\dots,i_{s}\}$. The right-hand side of (6.1) differs for $\pi^{\prime}$ and $\pi$ by the replacement of $h_{B}(\omega_{B})$ with
    $h_{B^{\prime}}(\omega_{B^{\prime}})=S_{n+1}(\omega_{B^{\prime}})S_{i_{s}}(\omega_{B^{\prime}})\dots S_{i_{2}}(\omega_{B^{\prime}})h_{i_{1}}(\omega_{B^{\prime}})=S_{n+1}(\omega_{B^{\prime}})h_{B}(\omega_{B^{\prime}}).$
    This agrees with the application of the operator $\mathfrak{m}(S_{n+1})$ to $T_{n}\dots T_{1}\xi$.
This completes the induction step and hence the proof. ∎
6.3. Fock space operators as limits of independent sums
Now we adapt Proposition 6.13 to the case of a sum of a creation, annihilation, and multiplication operators that will model limit distributions arising in applications of Theorem 1.1.
Proposition 6.14.
Let a Hilbert space $\mathcal{H}$ with unit vector $\xi$ be given, and let $X\in B(\mathcal{H})$ be self-adjoint. Let $\mathcal{H}^{\circ}=\mathcal{H}\ominus\mathbb{C}\xi$, and write $X$ in block form based on the decomposition $\mathcal{H}=\mathbb{C}\xi\oplus\mathcal{H}^{\circ}$ as
(6.2) $X=\begin{bmatrix}\alpha&h^{*}\\ h&S\end{bmatrix},\text{ where }\alpha\in\mathbb{C},h\in\mathcal{H}^{\circ},S\in B(\mathcal{H}^{\circ}).$
Fix a measure space $(\Omega,\rho)$ and nonnegative $w\in L^{\infty}(\Omega\times\Omega)$, and let $\mathcal{F}=\mathcal{F}(\Omega,\rho,w,\mathcal{H}^{\circ})$ be the associated Fock space. Let $A\subseteq\Omega$ with finite measure. Then define $Y\in B(\mathcal{F})$ by
$\widehat{X}=\alpha\mathfrak{n}(\mathbf{1}_{A})+\ell(\mathbf{1}_{A}h)+\ell(\mathbf{1}_{A}h)^{*}+\mathfrak{m}(\mathbf{1}_{A}S),$
where we view $\mathbf{1}_{A}h\in L^{2}(\Omega,\rho;\mathcal{H})$ and $\mathbf{1}_{A}S\in L^{\infty}(\Omega,B(\mathcal{H}^{\circ})$. Let $\mu$ and $\widehat{\mu}$ be the spectral distributions of $X$ and $\widehat{X}$ respectively with respect to the appropriate state vectors. Then
$m_{k}(\widehat{\mu})=\sum_{\pi\in\mathcal{NC}_{k}}\kappa_{\operatorname{bool},\pi}(\mu)\int_{A^{k}}\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{B},\omega_{\operatorname{pred}(B)})\prod_{B\in\pi}d\rho(\omega_{B}).$
Proof.
To compute $m_{k}(\widehat{\mu})=\langle\xi,\widehat{X}^{k}\xi\rangle_{\mathcal{F}}$, we expand $\widehat{X}^{k}=(\alpha\mathfrak{n}(\mathbf{1}_{A})+\ell(\mathbf{1}_{A}h)+\ell(\mathbf{1}_{A}h)^{*}+\mathfrak{m}(\mathbf{1}_{A}S))^{k}$ by multilinearity into the sum of $\langle\xi,T_{k}\dots T_{1}\xi\rangle_{\mathcal{F}}$, where $T_{j}\in\{\alpha\mathfrak{n}(\mathbf{1}_{A}),\ell(\mathbf{1}_{A}h),\ell(\mathbf{1}_{A}h)^{*},\mathfrak{m}(\mathbf{1}_{A}S)\}$. Then we apply Proposition 6.13 to each term. Each sequence of creation, annihilation, and multiplication operators such that $k(j)\geq 0$ has an associated non-crossing partition as in Observation 6.12. If the partition has unfinished blocks, then $T_{k}\dots T_{1}\xi$ is orthogonal to $\xi$ in $\mathcal{F}$ and hence $\langle\xi,T_{k}\dots T_{1}\xi\rangle$ vanishes. We are thus left with the terms where the partition does not have any unfinsished blocks. In this case, similar to Lemma 3.14, the partition $\pi$ uniquely determines the sequences of creation, annihilation, and multiplication operators by the rule that for singleton blocks $T_{j}=\alpha\mathfrak{n}(\mathbf{1}_{A})$, and for all other blocks, the leftmost element is the annihilation operator, the rightmost element is the creation operator, and the remaining terms are $\mathfrak{m}(\mathbf{1}_{A}S)$. Therefore, we obtain
(6.3) $\langle\xi,\widehat{X}^{k}\xi\rangle_{\mathcal{F}}=\sum_{\pi\in\mathcal{NC}_{k}}\int_{\Omega^{\times\pi}}\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{B},\omega_{\operatorname{pred}(B)})\prod_{B\in\pi}[\phi_{B}(\omega_{B})\,d\rho(\omega_{B})],$
where
$\phi_{B}(\omega)=\begin{cases}\alpha\mathbf{1}_{A}(\omega),&|B|=1\\ \mathbf{1}_{A}(\omega)\langle h,S^{|B|-2}h\rangle_{\mathcal{H}^{\circ}},&\text{else.}\end{cases}$
By Lemma 3.8 and equation (6.2), we have $\phi_{B}(\omega)=\mathbf{1}_{A}(\omega)\kappa_{\operatorname{bool},|B|}(\mu)$. Thus, (6.3) becomes
$\langle\xi,\widehat{X}^{k}\xi\rangle_{\mathcal{F}}=\sum_{\pi\in\mathcal{NC}_{k}}\prod_{B\in\pi}\kappa_{\operatorname{bool},|B|}(\mu)\int_{\Omega^{\times\pi}}\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{B},\omega_{\operatorname{pred}(B)})\prod_{B\in\pi}d\rho(\omega_{B}),$
which is the desired formula.o ∎
We now relate this back to the ideas of §5.1 by specializing to the case where $(\Omega,\rho)$ is a probability space and $w=\mathbf{1}_{\mathcal{E}}$ for some measurable $\mathcal{E}\subseteq\Omega\times\Omega$.
Corollary 6.15.
Consider the same setup as Proposition 6.14, and assume that $(\Omega,\rho)$ is a probability measure space, $\mathcal{E}\subseteq\Omega\times\Omega$ is measurable, and $w=\mathbf{1}_{E}$. Let $\widehat{\mu}$ be the distribution of the operator $\widehat{X}$ in that proposition. Then we have
(6.4) $m_{k}(\widehat{\mu})=\sum_{\pi\in NC(k)}\rho^{\times\pi}(\operatorname{Hom}(\mathrm{F}(\pi),(\Omega,\mathcal{E})))\kappa_{\operatorname{bool},\pi}(\mu).$
Now as in Proposition 1.2, let $G_{n}=(V_{n},E_{n})$ be a finite digraph for each $n\in\mathbb{N}$; let $(A_{n,v})_{v\in V_{n}}$ be a measurable partition of $\Omega$ into sets of measure $1/|V_{n}|$, and let ~$\tilde{\mathcal{E}}_{n}=\bigcup_{(v,w)\in E_{n}}A_{n,v}\times A_{n,w}$; suppose that $\rho(\tilde{E}_{n}\,\Delta\,\mathcal{E})\to 0$. If $\mu_{n}\in\mathcal{P}(\mathbb{R})$ such that $\mu_{n}^{\uplus|V_{n}|}\to\mu$, then $\boxplus_{G_{n}}(\mu_{n})\to\widehat{\mu}$.
Proof.
In Proposition 6.14, we take $A=\Omega$, and note that
$\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{\operatorname{pred}(B)},\omega_{B})=\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}\mathbf{1}_{\mathcal{E}}(\omega_{\operatorname{pred}(B)},\omega_{B})=\mathbf{1}_{\operatorname{Hom}(\mathrm{F}(\pi),(\Omega,E))}(\omega),$
and hence
$\int_{\Omega^{k}}\prod_{\begin{subarray}{c}B\in\pi\\ \operatorname{depth}(B)>1\end{subarray}}w(\omega_{\operatorname{pred}(B)},\omega_{B})\prod_{B\in\pi}d\rho(\omega_{B})=\rho^{\times\pi}(\operatorname{Hom}(\mathrm{F}(\pi),(\Omega,E))),$
so that
$m_{k}(\widehat{\mu})=\sum_{\pi\in NC(k)}\rho^{\times\pi}(\operatorname{Hom}(\mathrm{F}(\pi),(\Omega,\mathcal{E})))\kappa_{\operatorname{bool},\pi}(\mu).$
Now in the setting of Proposition 1.2, we have for rooted forests $G^{\prime}$ that
$\beta_{G^{\prime}}=\lim_{n\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\rho^{\times V^{\prime}}(\operatorname{Hom}(G^{\prime},(\Omega,\mathcal{E}))).$
Now from Lemma 4.3, it follows that
$\boxplus_{G_{n}}(\mu^{\uplus 1/|V_{n}|})\to\widehat{\mu}$
since the $k$th moments converge. From §4.2, it is clear that if $\mu_{n}^{\boxplus|V_{n}||}\to\mu$, then
$\lim_{n\to\infty}\boxplus_{G_{n}}(\mu_{n})=\lim_{n\to\infty}\boxplus_{G_{n}}(\mu^{\boxplus 1/|V_{n}|})=\widehat{\mu}.\qed$
Example 6.16 (BM Fock space and Brownian motion).
Fock spaces for BM independence associated to symmetric cones have been studied in [11]. As in §5.2, let $\Pi\subseteq\mathbb{R}^{d}$ be a closed convex salient cone. Consider the measure space $\Omega=\Pi$ with $\rho$ equal to the Lebesgue measure. For $\eta\in\Pi$, let
$\widehat{X}_{\eta}=\ell(\mathbf{1}_{[0,\eta]})+\ell(\mathbf{1}_{[0,\eta]})^{*}.$
Then $\widehat{X}_{\eta}$ is the BM Brownian motion. Note that if $\eta_{1}\preceq\eta_{2}$, then $\ell(\mathbf{1}_{[0,\eta_{1}]})+\ell(\mathbf{1}_{[0,\eta_{1}]})^{*}$ and $\ell(\mathbf{1}_{[\eta_{1},\eta_{2}]})+\ell(\mathbf{1}_{[\eta_{1},\eta_{2}]})^{*}$ are monotone independent. Similarly, if two intervals are elementwise incomparable, then the associated variables are boolean independent.
Example 6.17 (Multiregular digraphs).
As in §5.4, we consider multiregular digraphs $G_{n}=(V_{n},E_{n})$ with $V_{n}=\bigsqcup_{j=1}^{m}V_{n,j}$ so that |$|\{w\in V_{n,j}:v\rightsquigarrow w\}|=A_{n,i,j}$ for $v\in V_{n,i}$. Assume again that
$\lim_{n\to\infty}\frac{|V_{n,i}|}{|V_{n}|}=t_{i},\qquad\lim_{n\to\infty}\frac{A_{n,i,j}}{|V_{n}|}=a_{i,j},$
and recall by Proposition 5.11 that for a rooted tree $G^{\prime}$,
$\lim_{n\to\infty}\frac{|\operatorname{Hom}(G^{\prime},G_{n})|}{|V_{n}|^{|V^{\prime}|}}=\sum_{\ell:V^{\prime}\to[m]}t_{\ell(r)}\prod_{v\in V^{\prime}\setminus\{r\}}a_{\ell(v_{-}),\ell(v)}=:\beta_{G^{\prime}}.$
These coefficients can be realized with a tuple $(\Omega,\rho,w)$ as follows. Let
$\Omega=[m],\qquad\rho=\sum_{i=1}^{m}t_{i}\delta_{i},\qquad w(i,j)=\frac{a_{i,j}}{t_{j}}.$
Then a direct computation shows that for a rooted tree $G^{\prime}=(V^{\prime},E^{\prime})$,
$\int_{\Omega^{V^{\prime}}}\prod_{v\in V^{\prime}\setminus\{r\}}w(\omega_{v_{-}},\omega_{v})\,d\rho^{\times V^{\prime}}(\omega)=\sum_{\ell:V^{\prime}\to[m]}t_{\ell(r)}\prod_{v\in V^{\prime}\setminus\{r\}}a_{\ell(v_{-}),\ell(v)}.$
Therefore, the construction in Proposition 6.14 with this choice of $(\Omega,\rho,w)$ will realize the moments of limit distributions from Proposition 5.11.